### v2

# bibliography.bib

Abstract
Consider a finite dimensional real vector space and a finite group acting unitarily on it. We study the general problem of constructing Euclidean stable embeddings of the quotient space of orbits. Our embedding is based on subsets of sorted coorbits. Our main result shows that, whenever such embeddings are injective, they are automatically bi-Lipschitz. Additionally, we demonstrate that stable embeddings can be achieved with reduced dimensionality, and that any continuous or Lipschitz $G$-invariant map can be factorized through these embeddings.
1 Introduction
In a lot of machine learning problems we want to embed our data into an Euclidean space $\mathbb{R}^{m}$ using a symmetry-invariant embedding $\Psi$ and utilize $\mathbb{R}^{m}$ as our feature space. This embedding $\Psi$ should also separate data orbits and satisfy certain stability conditions to ensure that small perturbations of the input do not significantly impact the predictions. We worked in the orbit separation problem in [balan2023inj]. In this paper we focus on the stability problem.
This problem is an instance of invariant machine learning. [mixon2022injectivity, aslan2023group, dym2022low, bronstein, maron2018invariant, DUFRESNE20091979, cahill2023bilipschitz].
The most common group actions in invariant machine learning are permutations [sannai2020universal, cahill22, balan2022permutation] reflections [mixon2022max] and translations [Cahill19]. A related, by somewhat different problem is the case of equivariant embeddings [lipman2022, maron2018invariant, villar2021scalars, villar2022dimensionless, blum2022equivariant].
Our work extends and unifies two prior approaches: the max filter embedding introduced in [cahill22, mixon2022injectivity], and the permutation invariant representation introduced in [balan2022permutation].
The construction of permutation invariant embeddings is closely connected to the phase retrieval problem [balan2004signal, bcmn] that has a large body of results. For instance, [balan16] provides exact estimates for both the upper and lower Lipschitz bounds, addressing both the real and complex cases of the phase retrieval problem. In [Balan2023Rel] we establish an isometric identification of the real phase retrieval problem to $S_{2}$-invariant representations.
A completely different approach is considered in [eriksson2018quantitative]. There it is proved that for any discrete group $G$ of isometries acting on $\mathbb{R}^{d}$, it is possible to construct a bi-Lipschitz map from $\mathbb{R}^{d}/G$ to $\mathbb{R}^{N}$, where the distortion depends solely on $d$.
Since the first draft of this paper was placed on arxiv, the authors of [qaddura2024stable] extended the construction of this paper and proved that for certain compact groups, given enough generic templates, the coorbit filter bank (as defined there) is injective and bi-Lipschitz.
In this paper we construct an Euclidean embedding that is globally bi-Lipschitz and can be implemented relatively easy. Our paper is organized as follows: Chapter 1 introduces the embedding map. Chapter 2 explores the upper Lipschitz bound for the proposed embeddings and establishes the equivalence between injectivity on the quotient space and stability. Chapter 3 demonstrates that a generic linear projection can be used to reduce the dimension of the target space while preserving both injectivity and stability. Finally, chapter 4 presents universal factorization results for arbitrary continuous or Lipschitz $G$-invariant maps.
1.1 Notation
In this paper we use the same notation as in [balan2023inj]. Let $({\mathscr{V}},{\langle\cdot,\cdot\rangle})$ be a $d$-dimensional real vector space with scalar product, and $d\geq 2$. Assume $(G,\cdot)$ is a finite group of order $|G|=N$ acting unitarily on ${\mathscr{V}}$. For every $g\in G$, we denote by $U_{g}x$ the group action on vector $x\in{\mathscr{V}}$. Let $\hat{{\mathscr{V}}}={\mathscr{V}}/\sim$ denote the quotient space with respect to this group action, $x\sim y$ if and only if $y=U_{g}x$ for some $g\in G$. We denote by $[x]$ the orbit of vector $x$, i.e. $[x]=\{U_{g}x:g\in G\}$. The natural metric, $\operatorname{{\mathbf{d}}}:\hat{{\mathscr{V}}}\times\hat{{\mathscr{V}}}\to\mathbb{R}$, is defined by
$\operatorname{{\mathbf{d}}}([x],[y])=\min_{h_{1},h_{2}\in G}\lVert U_{h_{1}}x-U_{h_{2}}y\rVert=\min_{g\in G}\lVert x-U_{g}y\rVert.$ (1)
Note $(\hat{{\mathscr{V}}},\operatorname{{\mathbf{d}}})$ is a complete metric space.
Our goal is to construct a bi-Lipschitz Euclidean embedding of the metric space $(\hat{{\mathscr{V}}},\operatorname{{\mathbf{d}}})$ into an Euclidean space $\mathbb{R}^{m}$.
Specifically, we want to construct a function $\Psi:{\mathscr{V}}\to\mathbb{R}^{m}$ such that
  1. 1.
    $\Psi(U_{g}x)=\Psi(x),\ \forall x\in{\mathscr{V}},\ \forall g\in G$,
  2. 2.
    If $x,y\in{\mathscr{V}}$ are such that $\Psi(x)=\Psi(y)$, then there exist $g\in G$ such that $y=U_{g}x$,
  3. 3.
    There are $0<a<b<\infty$ such that for any $x,y\in{\mathscr{V}}$
    $a\operatorname{{\mathbf{d}}}([x],[y])^{2}\leq\lVert\Psi(x)-\Psi(y)\rVert^{2}\leq b(\operatorname{{\mathbf{d}}}([x],[y]))^{2}.$
The invariance property (1) allows to lift $\Psi$ to a map $\hat{\Psi}$ acting on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/\sim$:
$\hat{\Psi}:\hat{{\mathscr{V}}}\rightarrow\mathbb{R}^{m},\quad\hat{\Psi}([x])=\Psi(x),\quad\forall[x]\in\hat{{\mathscr{V}}}.$
If a $G$-invariant map $\Psi$ satisfies property (2) we say that $\Psi$ separates $G$-orbits in ${\mathscr{V}}$.
[…]
Notation 1.1.
$\operatorname{\downarrow}:\mathbb{R}^{r}\to\mathbb{R}^{r}$ denoted the operator that takes as input a vector in $\mathbb{R}^{r}$ and returns a monotonically decreasing sorted vector of same length $r$ that has same entries as the input vector:
$x\in\mathbb{R}^{r}\mapsto\operatorname{\downarrow}x=(x_{\sigma(i)})_{1\leq i\leq r}\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ x_{\sigma(1)}\geq\cdots\geq x_{\sigma(r)}$
for some permutation $\sigma:\{1,\ldots,r\}\rightarrow\{1,\ldots,r\}$.
For an integer $p\in\mathbb{N}$, we denote $[p]=\{1,2,\ldots,p\}$. For a set $S$, $|S|$ denotes its cardinal. Fix a $p$-tuple of vectors ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$. For any $i\in[p]$ and $j\in[N]$ we define the operator $\Phi_{w_{i},j}:{\mathscr{V}}\to\mathbb{R}$ so that $\Phi_{w_{i},j}(x)$ is the $j$-th coordinate of the sorted vector $\operatorname{\downarrow}({\langle U_{g}w_{i},x\rangle})_{g\in G}$. Fix a set $S\subset[N]\times[p]$ such that $|S|=m$, and for $i\in[p]$, let $S_{i}=\{k\in[N]:(k,i)\in S\}$ (the $i^{th}$ column of $S$). Denote by $m_{i}$ the cardinal of the set $S_{i}$, $m_{i}=|S_{i}|$. Thus $m=\sum_{i=1}^{p}m_{i}$.
Notation 1.2.
The coorbit embedding $\Phi_{{\textbf{w}},S}$ associated to windows ${\textbf{w}}\in{\mathscr{V}}^{p}$ and index set $S\subset[N]\times[p]$ is given by the map
$\Phi_{{\textbf{w}},S}:{\mathscr{V}}\rightarrow\mathbb{R}^{m}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \Phi_{{\textbf{w}},S}(x)=[\{\Phi_{w_{1},j}(x)\}_{j\in S_{1}},\dots,\{\Phi_{w_{p},j}(x)\}_{j\in S_{p}}]\in\mathbb{R}^{m}.$ (2)
$[y]$$[x]$$\hat{{\mathscr{V}}}$$\mathbb{R}^{m}$$\mathbb{R}^{q}$$\Phi_{{\textbf{w}},S}$$\Phi_{{\textbf{w}},S}$$\ell$$\ell$$\Phi_{{\textbf{w}},S}([x])$$\Phi_{{\textbf{w}},S}([y])$$\ell\circ\Phi_{{\textbf{w}},S}([y])$$\ell\circ\Phi_{{\textbf{w}},S}([x])$
Figure 1: Proposed embedding scheme
Let $\ell:\mathbb{R}^{m}\to\mathbb{R}^{q}$ be a linear map.
Notation 1.3.
The embedding $\Psi_{{\textbf{w}},S,\ell}$ associated to windows ${\textbf{w}}\in{\mathscr{V}}^{p}$, index set $S\subset[N]\times[p]$ and linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{q}$ is given by the map
$\Psi_{{\textbf{w}},S,\ell}=\ell\circ\Phi_{{\textbf{w}},S}:{\mathscr{V}}\to\mathbb{R}^{q}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \Psi_{{\textbf{w}},S,\ell}(x)=\ell(\Phi_{{\textbf{w}},S}(x))$ (3)
obtained by composition of $\ell$ with the coorbit embedding $\Phi_{{\textbf{w}},S}$.
In this paper we focus on stability and universality properties of maps $\Phi_{{\textbf{w}},S}$ and $\Psi_{{\textbf{w}},S,\ell}$.
Informally, our main results, 1.4, 1.6, 1.7 and 1.8, state that: (1) "injectivity" implies "(bi-Lipschitz) stability", (2) stable bi-Lipschitz embedding can be achieved into an Euclidean space of dimension at most twice the dimension of the input space; and (3) any continuous or even Lipschitz map factors through $\Phi_{{\textbf{w}},S}$.
For the rest of the paper we shall use interchangeably $\Phi_{i,j}$ instead of $\Phi_{w_{i},j}$.
1.2 Main results
In this section, we summarize the key results of this paper, that focus on bi-Lipschitz stability of embeddings, dimensionality reduction, and universality of $G$-invariant maps.
Theorem 1.4.
Let $G$ be a finite group acting unitarily on the vector space ${\mathscr{V}}$. For fixed ${\textbf{w}}\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$, where $|S|=m$, suppose that the map $\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\to\mathbb{R}^{m}$ is injective on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$. Then, there exist $0<a\leq b<\infty$ such that for all $x,y\in{\mathscr{V}}$,
$a\operatorname{{\mathbf{d}}}([x],[y])\leq\lVert\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)\rVert_{2}\leq b\operatorname{{\mathbf{d}}}([x],[y]).$
Corollary 1.5.
When injective, the Max Filter embedding is bi-Lipschitz. Specifically: Let $G$ be a finite group acting unitarily on $V$. Suppose the max filter bank embedding $\hat{\Phi}_{w,S_{\text{max}}}:\hat{{\mathscr{V}}}\to\mathbb{R}^{p}$, defined by
$\hat{\Phi}_{w,S_{\text{max}}}([x])=\left(\max_{g\in G}\langle x,g\cdot w_{k}\rangle\right)_{k\in[p]}$
is injective. Then $\hat{\Phi}_{w,S_{\text{max}}}$ is bi-Lipschitz, that is, there are $0<a\leq b<\infty$, so that for every $x,y\in{\mathscr{V}}$,
$a\cdot\operatorname{{\mathbf{d}}}([x],[y])\leq\|\Phi_{w,S_{\text{max}}}(x)-\Phi_{w,S_{\text{max}}}(y)\|_{2}\leq b\cdot\operatorname{{\mathbf{d}}}([x],[y]).$
Theorem 1.6.
Let $\{U_{g}\leavevmode\nobreak\ ,\leavevmode\nobreak\ g\in G\}$ denote a representation of a finite group $G$ of order $N$ acting by isometries on the real vector space ${\mathscr{V}}$ of dimension $d$. Let $V_{G}=\{x\in{\mathscr{V}}:U_{g}x=x\leavevmode\nobreak\ ,\leavevmode\nobreak\ \forall g\in G\}$ denote the linear space of vectors invariant of this representation, and let $d_{G}=\dim(V_{G})$ denote its dimension. Let ${\textbf{w}}\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$ so that $\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\to\mathbb{R}^{m}$ is injective on the quotient space $\hat{{\mathscr{V}}}$. Then, for a generic linear map $\ell:\mathbb{R}^{m}\to\mathbb{R}^{2d-d_{G}}$, the map $\hat{\Psi}_{{\textbf{w}},S,\ell}=\ell\circ\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\rightarrow\mathbb{R}^{2d-d_{G}}$ is injective and bi-Lipschitz. Here generic means open dense with respect to Zariski topology over the set of matrices.
Theorem 1.7.
Assume that for some $S\in[N]\times[p]$ and ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$, the map $\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\to\mathbb{R}^{m}$ is bi-Lipschitz on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$, with (upper) Lipschitz constant $b=Lip(\hat{\Phi}_{{\textbf{w}},S})$ and lower Lipschitz constant $a>0$.
  1. 1.
    For every Lipschitz map $F:\hat{{\mathscr{V}}}\to H$ into a real Hilbert space $H$ there exists a Lipschitz map $T:\mathbb{R}^{m}\rightarrow H$ such that $F=T\circ\Phi_{{\textbf{w}},S}$ and has Lipschitz constant $\operatorname{Lip}(T)\leq\frac{1}{a}\operatorname{Lip}(F)$.
  2. 2.
    Conversely, for any Lipschitz map $T:\mathbb{R}^{m}\to H$, the map $F=T\circ\Phi_{{\textbf{w}},S}$ is Lipschitz, with Lipschitz constant $\operatorname{Lip}(F)\leq b\operatorname{Lip}(T)$.
Theorem 1.8.
Assume that, for fixed $S\in[N]\times[p]$ and ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$, the map $\hat{\Phi}_{{\textbf{w}},S}:{\mathscr{V}}\to\mathbb{R}^{m}$ is bi-Lipschitz on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$. Let $L$ denote a locally convex topological vector space. For every continuous map $F:\hat{{\mathscr{V}}}\to L$ there exists a continuous $T:\mathbb{R}^{m}\to L$ such that $F=T\circ\Phi_{{\textbf{w}},S}$. Furthermore $T(\mathbb{R}^{m})$ is included in the convex hull of $F(\hat{{\mathscr{V}}})$.
2 Stability of Embedding
Suppose that for ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$ the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective. In this case, we claim the map $\hat{\Phi}_{{\textbf{w}},S}$ is also bi-Lipschitz, (1.4). An estimate of the upper Lipschitz constant $b$ is given in 2.1.
1.5 follows from 1.4.
The proof of 1.4 is contained in the next two subsections.
2.1 Upper Lipschitz bound
Note that $\Phi_{{\mathbf{w}},S}$ is Lipschitz because it is a composition of Lipschitz maps. It remains to estimate the upper bound.
Lemma 2.1.
Consider $G$ be a finite group of size $N$ acting unitarily on ${\mathscr{V}}$. Let ${\mathbf{w}}\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$. Denote
$B=\max_{\begin{subarray}{c}\sigma_{1},\ldots,\sigma_{p}\subset G\\ |\sigma_{i}|=m_{i},\forall i\end{subarray}}\lambda_{max}\left(\sum_{i=1}^{p}\sum_{g\in\sigma_{i}}U_{g}w_{i}w_{i}^{T}U_{g}^{T}\right),$
where $S_{i}=\{j\in[N],(i,j)\in S\}$ and $m_{i}=|S_{i}|$. Then $\hat{\Phi}_{{\mathbf{w}},S}:({\widehat{V}},\operatorname{{\mathbf{d}}})\rightarrow\mathbb{R}^{m}$ is Lipschitz with an upper Lipschitz constant $\sqrt{B}$.
Proof.
For fixed $x,y\in{\mathscr{V}}$, $i\in[p],j\in S_{i}$, let $\psi_{i,j}:[0,1]\to\mathbb{R}$, where
$\psi_{i,j}(t)=\Phi_{i,j}((1-t)x+ty)={\langle(1-t)x+ty,U_{g(t)}w_{i}\rangle}.$
By Lebesgue differentiation theorem we have that $\psi_{i,j}$ is differentiable almost everywhere. Consequently $\Phi_{i,j}$ is also differentiable almost everywhere. Notice that
$\frac{d}{dt}\psi_{i,j}(t)={\langle y-x,U_{g(t)}w_{i}\rangle}$
for almost every $t$. Specifically, $\psi_{i,j}(t)$ is differentiable at all $t\in[0,1]$ such that there exists $\epsilon>0$ so that $g{|}_{(t-\epsilon,t+\epsilon)}$ can be chosen to be constant. This happens because $G$ is finite.
By the fundamental theorem of calculus we get
[…]
Therefore,
[…]
so
[…]
$\leq\sqrt{B_{i}}\lVert x-y\rVert$
where $B_{i}=\max_{\begin{subarray}{c}\sigma\subset G,\\ |\sigma|=|S_{i}|,\end{subarray}}\lambda_{max}\left(\sum_{g\in\sigma}U_{g}w_{i}w_{i}^{T}U_{g}^{T}\right).$
Hence, for $x\in[x],y\in[y]$ so that $\operatorname{{\mathbf{d}}}([x],[y])=\lVert x-y\rVert$,
$\lVert\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)\rVert^{2}=\sum_{i=1}^{p}\lVert\Phi_{i,j}(y)-\Phi_{i,j}(x)\rVert^{2}$ $\leq\sum_{i=1}^{p}B_{k}\lVert x-y\rVert^{2}$
$\leq B\operatorname{{\mathbf{d}}}([x],[y])^{2}.$
2.2 Lower Lipschitz bound
We start with some useful geometric results.
2.2.1 Geometric Analysis of Coorbits
Fist let us introduce some additional notation. For fixed $i\in[p]$, $j\in[N]$ and $x\in{\mathscr{V}}$ we define the following non-empty subset of the group $G$:
$L^{i,j}(x)=\{g\in G:{\langle x,U_{g}w_{i}\rangle}=\Phi_{i,j}(x)\}.$ (4)
This represents the collection of group elements that achieve the $j$-th position for the sorted co-orbit $\downarrow({\langle U_{g}w_{i},x\rangle})_{g\in G}$.
Consider also the map
$\Delta^{i,j}(x)=\begin{cases}\min_{g\notin L^{i,j}(x)}(|{\langle U_{g}w_{i},x\rangle}-\Phi_{i,j}(x)|)\frac{1}{\lVert w_{i}\rVert},\leavevmode\nobreak\ \text{ if }L^{i,j}(x)\not=G\\ \frac{\lVert x\rVert}{\lVert w_{i}\rVert},\leavevmode\nobreak\ \text{ if }L^{i,j}(x)=G.\\ \end{cases}$ (5)
Lemma 2.2.
a. For any $x\in{\mathscr{V}}$, $i\in[p]$, and $j\in[N]$,
|$\left|\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}>\Phi_{i,j}(x)\}\right|\leq j-1$ (6)
|$\left|\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)\}\right|\leq N-j.$ (7)
b. For any $x\in{\mathscr{V}}$, $i\in[p]$, and $j\in[N-1]$,
(i) either $\Phi_{i,j}(x)=\Phi_{i,j+1}(x)$, in which case $L^{i,j}(x)=L^{i,j+1}(x)$,
or
(ii) $\Phi_{i,j}(x)>\Phi_{i,j+1}(x)$, in which case $L^{i,j}(x)\neq L^{i,j+1}(x)$ and
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}>\Phi_{i,j+1}(x)\}=\cup_{k\leq j}L^{i,k}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \left|\cup_{k\leq j}L^{i,k}(x)\right|=j$
and
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)\}=\cup_{k\geq j+1}L^{i,k}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \left|\cup_{k\geq j+1}L^{i,k}(x)\right|=N-j.$
Proof.
(a) Recall that $\Phi_{i,j}(x)$, is the $j$-th coordinate of the monotonically decreasing sorted vector $\operatorname{\downarrow}({\langle U_{g}w_{i},x\rangle})_{g\in G}$. Suppose that
|$\left|\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}>\Phi_{i,j}(x)\}\right|>j-1.$
Then there are at least $j$, distinct elements of group $G$, $(h_{1},\dots h_{j})$, such that ${\langle U_{h_{k}}w_{i},x\rangle}>{\langle U_{g}w_{i},x\rangle},\leavevmode\nobreak\ \forall k\in[j]$. But this is a contradiction. Similarly, if
|$\left|\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)\}\right|>N-j$
then there exist at least $N-j+1$, distinct elements of group $G$, say $\{h_{1},\dots h_{j}\}$ such that ${\langle U_{h_{k}}w_{i},x\rangle}<{\langle U_{g}w_{i},x\rangle}\leavevmode\nobreak\ \forall k\in[N-j+1]$, which is also a contradiction.
(b) If $\Phi_{i,j}(x)=\Phi_{i,j+1}(x)$, a group element $g$ achieves $\Phi_{i,j}(x)$ if and only if, it also achieves $\Phi_{i,j+1}(x)$, therefore $L^{i,j}(x)=L^{i,j+1}(x)$. On the other hand, if $\Phi_{i,j}(x)>\Phi_{i,j+1}(x)$ then we claim that $L^{i,j}(x)$ and $L^{i,j+1}(x)$ are disjoint sets, because otherwise there is $g\in L^{i,j}(x)\cap L^{i,j+1}(x)$, but then $\Phi_{i,j}(x)={\langle U_{g}w_{i},x\rangle}=\Phi_{i,j+1}(x)$.
Now assume that
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}>\Phi_{i,j+1}(x)\}\not=j.$
Without loss of generality
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}>\Phi_{i,j+1}(x)\}>j$
so there exists at least $j+1$ group elements $(h_{1},\dots,h_{j+1})$, such that
${\langle U_{h_{k}}w_{i},x\rangle}>\Phi_{i,j+1},\leavevmode\nobreak\ \forall k\in[j+1]$
but this is a contradiction. Similarly,
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)\}=\cup_{k\geq j+1}L^{i,k}(x)\leavevmode\nobreak\ \leavevmode\nobreak\ ,\leavevmode\nobreak\ \leavevmode\nobreak\ \left|\cup_{k\geq j+1}L^{i,k}(x)\right|=N-j.$
Note that for any $w_{1},\ldots,w_{p}\in{\mathscr{V}}\setminus\{0\}$ the subset $L^{i,j}(x)\subset G$ has the following “nestedness” property.
Lemma 2.3.
For any $x,y\in{\mathscr{V}}$ such that $\lVert y\rVert<\frac{1}{2}\Delta^{i,j}(x)$, we have that $L^{i,j}(x+y)\subset L^{i,j}(x)$. Furthermore,
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}>\Phi_{i,j}(x)\}\subset\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x+y\rangle}>\Phi_{i,j}(x+y)\},$
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)\}\subset\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x+y\rangle}<\Phi_{i,j}(x+y)\},$
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x+y\rangle}\geq\Phi_{i,j}(x+y)\}\subset\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}\geq\Phi_{i,j}(x)\}$
and
$\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x+y\rangle}\leq\Phi_{i,j}(x+y)\}\subset\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}\leq\Phi_{i,j}(x)\}.$
Proof.
Suppose that exists $g\in G$ such that $g\in L^{i,j}(x+y)$ but $g\notin L^{i,j}(x)$. Without loss of generality assume that ${\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)$. Then for every $h\in\cup_{k\leq j}L^{i,k}(x)$
${\langle U_{h}w_{i},x+y\rangle}-{\langle U_{g}w_{i},x+y\rangle}\geq{\langle U_{h}w_{i},x\rangle}-{\langle U_{g}w_{i},x\rangle}-2\lVert y\rVert\lVert w_{i}\rVert>0.$
On the other hand, ${\langle U_{g}w_{i},x+y\rangle}=\Phi_{i,j}(x+y)$. Thus
$\cup_{k\leq j}L^{i,k}(x)\subset\{h\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{h}w_{i},x+y\rangle}>\Phi_{i,j}(x+y)\}.$
But the set $\cup_{k\leq j}L^{i,k}(x)$ contains at least $j$ elements (since each $L^{i,j}(x)$ is non-empty) and so we derived a contradiction with 2.2(a) Equation 6.
Lemma 2.4.
For $i\in[p]$ and $j\in[N]$, fix vectors $x,y\in{\mathscr{V}}$ and positive numbers $c_{1},c_{2}>0$ such that $\max(c_{1},c_{2})\lVert y\rVert<\frac{1}{4}\Delta^{i,j}(x)$. Then $L^{i,j}(x+c_{1}y)=L^{i,j}(x+c_{2}y)$.
Proof.
Assume that exist $g_{1}\in L^{i,j}(x+c_{2}y)$ with $g_{1}\notin L^{i,j}(x+c_{1}y)$. Without loss of generality assume that ${\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$. Let $q>j$ be the smallest integer such that $g_{1}\in L^{i,q}(x+c_{1}y)$. Then $\Phi_{i,q}(x+c_{1}y)={\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$. By 2.2 (b)(ii),
$\left|\cup_{r\leq j}L^{i,r}(x+c_{1}y)\right|=q-1\geq j,$
and $g_{1}\not\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$. On the other hand, from 2.2 (a), Equation 6,
|$\left|\{h\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{h}w_{i},x+c_{2}y\rangle}>\Phi_{i,j}(x+c_{2}y)\}\right|\leq j-1$
Hence
$\cup_{r\leq j}L^{i,r}(x+c_{1}y)\setminus\{h\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{h}w_{i},x+c_{2}y\rangle}>\Phi_{i,j}(x+c_{2}y)\}\neq\emptyset$
Therefore there exists $h\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$ such that
${\langle U_{h}w_{i},x+c_{2}y\rangle}\leq\Phi_{i,j}(x+c_{2}y)={\langle U_{g_{1}}w_{i},x+c_{2}y\rangle}.$ (8)
On the other hand, by 2.3, $g_{1}\in L^{i,j}(x)$.
But if ${\langle U_{h}w_{i},x\rangle}-{\langle U_{g_{1}}w_{i},x\rangle}>0$ then
${\langle U_{h}w_{i},x+c_{2}y\rangle}-{\langle U_{g_{1}}w_{i},x+c_{2}y\rangle}\geq\lVert w_{i}\rVert(\Delta^{i,j}(x)-2c_{2}\lVert y\rVert)>0$
which is a contradiction with (8). If ${\langle U_{h}w_{i},x\rangle}-{\langle U_{g_{1}}w_{i},x\rangle}<0$ then
${\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}-{\langle U_{h}w_{i},x+c_{1}y\rangle}\geq\lVert w_{i}\rVert(\Delta^{i,j}(x)-2c_{1}\lVert y\rVert)>0$
which is a contradiction with $h\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$. Therefore ${\langle U_{h}w_{i},x\rangle}={\langle U_{g_{1}}w_{i},x\rangle}$ and thus $h\in L^{i,j}(x)$. But then
$0\geq{\langle U_{h}w_{i},x+c_{2}y\rangle}-{\langle U_{g_{1}}w_{i},x+c_{2}y\rangle}$ $={\langle U_{h}w_{i},c_{2}y\rangle}-{\langle U_{g_{1}}w_{i},c_{2}y\rangle}$
$=c_{2}({\langle U_{h}w_{i},y\rangle}-{\langle U_{g_{1}}w_{i},y\rangle})$
and
$0<{\langle U_{h}w_{i},x+c_{1}y\rangle}-{\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}$ $={\langle U_{h}w_{i},c_{1}y\rangle}-{\langle U_{g_{1}}w_{i},c_{1}y\rangle}$
$=c_{1}({\langle U_{h}w_{i},y\rangle}-{\langle U_{g_{1}}w_{i},y\rangle}).$
Which contradict each other since $c_{1},c_{2}>0$.
Lemma 2.5.
For any $w_{1},\ldots,w_{p}\in{\mathscr{V}}\setminus\{0\}$ and $x\in{\mathscr{V}}$, the sets $L^{i,j}(x)$ and perturbation bounds $\Delta^{i,j}(x)$ have the following properties:
  1. 1.
    For any $t>0$, $L^{i,j}(tx)=L^{i,j}(x)$.
  2. 2.
    For any $i\in[p]$, $j\in[N]$, and $t>0$, $\Delta^{i,j}(tx)=t\Delta^{i,j}(x)$.
  3. 3.
    For any $i\in[p]$, $j\in[N]$, and $x\in{\mathscr{V}}\setminus\{0\}$, $\Delta^{i,j}(x)>0$.
Proof.
1.,2. For $t>0$, $\Phi_{i,j}(tx)=t\Phi_{i,j}(x)$, from where the claims follow from the definitions of $L^{i,j}(x)$ and $\Delta^{i,j}(x)$.
3. This claim follows from definitions of $\Delta_{i,j}$ which is the minimum of a finite set of positive numbers.
Lemma 2.6.
Fix $w_{i}\in{\mathscr{V}}\setminus\{0\}$ and $j\in[N]$. For any $k>1$, fix $z_{1}\in{\mathscr{V}}$ of unit norm, $\lVert z_{1}\rVert=1$, and choose arbitrary $z_{2},...,z_{k}$ that satisfy only the norm conditions
$\lVert z_{l+1}\rVert\leq\min(\frac{1}{4}\Delta^{i,j}(\sum_{r=1}^{l}z_{r}),\frac{1}{4}\lVert z_{l}\rVert)\leavevmode\nobreak\ \forall l\in[k-1].$
For any scalars $a_{1},\dots,a_{k}\in\left(1-\frac{1}{16k}\Delta^{i,j}(\sum_{r=1}^{k}z_{r}),1+\frac{1}{16k}\Delta^{i,j}(\sum_{r=1}^{k}z_{r})\right)$ the following hold true:
  1. 1.
    If $g_{1},g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then
    ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}.$
  2. 2.
    $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})=L^{i,j}(\sum_{r=1}^{k}z_{r}),$ (9)
  3. 3.
    $\frac{1}{4}\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})<\Delta^{i,j}(\sum_{r=1}^{k}z_{r})<4\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}).$ (10)
  4. 4.
    $\bigcup_{l\leq j}L^{i,l}(\sum_{r=1}^{k}a_{r}z_{r})=\bigcup_{l\leq j}L^{i,l}(\sum_{r=1}^{k}z_{r}),$ (11)
    $\bigcup_{l\geq j}L^{i,l}(\sum_{r=1}^{k}a_{r}z_{r})=\bigcup_{l\geq j}L^{i,l}(\sum_{r=1}^{k}z_{r}).$ (12)
  5. 5.
    For every $e\in{\mathscr{V}}$ with $\lVert e\rVert<\frac{1}{16}\Delta^{i,j}(\sum_{r=1}^{k}z_{r})$,
    $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)=L^{i,j}(\sum_{r=1}^{k}z_{r}+e)$ (13)
Remark 2.7.
Notice this Lemma allows us to choose arbitrary directions for vectors $z_{2},...,z_{k}$. Conclusions of the lemma hold true even for $z_{2}=\ldots=z_{k}=0$. On the other hand, the norm conditions $\lVert z_{l+1}\rVert\leq\frac{1}{4}\lVert z_{l}\rVert$ and $\lVert z_{1}\rVert=1$ prevent $\sum_{l=1}^{k}z_{l}$ from ever reaching 0. Hence $\Delta({i,j}(\sum_{r=1}^{l}z_{r})>0$ for all $l\in[k]$.
Proof.
1. Note that by norm conditions on $z_{r}$, 2.3 and 2.4, if $g_{1},g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}\leavevmode\nobreak\ \forall l\in[k]$. Starting with $l=1$ and proceeding recursively we get ${\langle U_{g_{1}}w_{i},z_{r}\rangle}={\langle U_{g_{2}}w_{i},z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$. Therefore ${\langle U_{g_{1}}w_{i},a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},a_{r}z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$ and so ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}.$
2.
$\subset$". From 2.3 we have that $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$.
$\supset$". Let $g_{1}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$. Take any $g_{2}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. By 2.3, $g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$, but from part 1 of this Lemma,
${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}.$
Therefore $g_{1}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$, which proves the other inclusion.
3.
Case 1. Suppose $L^{i,j}(\sum_{r=1}^{k}z_{r})=G$. From 2.3 we know that
$G=L^{i,j}(\sum_{r=1}^{k}z_{r})\subset L^{i,j}(\sum_{r=1}^{k-1}z_{r})\subset\dots\subset L^{i,j}(z_{1}).$
herefore, $L^{i,j}(z_{r})=G,\leavevmode\nobreak\ \forall r\in[k]$ and consequently $L^{i,j}(a_{r}z_{r})=G,\leavevmode\nobreak\ \forall r\in[k]$.
Moreover, $a_{1},\dots a_{k}\in(7/8,9/8)$. Therefore,
$\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})=\frac{1}{\lVert w_{i}\rVert}\lVert\sum_{r=1}^{k}a_{r}z_{r}\rVert\leq\frac{9}{8\lVert w_{i}\rVert}\left(\sum_{r=1}^{k}\lVert z_{r}\rVert\right)<\frac{3}{2\lVert w_{i}\rVert}\lVert z_{1}\rVert$
and
$\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})=\frac{1}{\lVert w_{i}\rVert}\lVert\sum_{r=1}^{k}a_{r}z_{r}\rVert\geq\frac{1}{\lVert w_{i}\rVert}\left(\frac{7}{8}\lVert z_{1}\rVert-\frac{9}{8}\sum_{r=2}^{k}\lVert z_{r}\rVert\right)>\frac{1}{2\lVert w_{i}\rVert}\lVert z_{1}\rVert.$
Similarly,
$\Delta^{i,j}(\sum_{r=1}^{k}z_{r})=\frac{1}{\lVert w_{i}\rVert}\lVert\sum_{r=1}^{k}z_{r}\rVert\leq\frac{1}{\lVert w_{i}\rVert}\left(\sum_{r=1}^{k}\lVert z_{r}\rVert\right)<\frac{4}{3\lVert w_{i}\rVert}\lVert z_{1}\rVert$
and
$\Delta^{i,j}(\sum_{r=1}^{k}z_{r})=\frac{1}{\lVert w_{i}\rVert}\lVert\sum_{r=1}^{k}z_{r}\rVert\geq\frac{1}{\lVert w_{i}\rVert}\left(\lVert z_{1}\rVert-\sum_{r=2}^{k}\lVert z_{r}\rVert\right)>\frac{2}{3\lVert w_{i}\rVert}\lVert z_{1}\rVert.$
So,
$\frac{1}{4}\Delta^{i,j}(\sum_{r=1}^{k}z_{r})\leq\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\leq 4\Delta^{i,j}(\sum_{r=1}^{k}z_{r}).$
Case 2. Now assume that $L^{i,j}(\sum_{r=1}^{k}z_{r})\not=G$. Fix $g_{1}\in G$ that achieves $\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$, i.e.
$\frac{1}{\lVert w_{i}\rVert}|{\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}-\Phi_{i,j}(\sum_{r=1}^{k}a_{r}z_{r})|=\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$
and $g_{2}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$.
Then
$\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})=$ $\frac{1}{\lVert w_{i}\rVert}\left|{\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}-{\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}\right|$
$\geq$ $\frac{1}{\lVert w_{i}\rVert}\left|{\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}z_{r}\rangle}-{\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}z_{r}\rangle}\right|-\sum_{r=1}^{k}2|1-a_{r}|\lVert z_{r}\rVert$
$\geq$ $\Delta^{i,j}(\sum_{r=1}^{k}z_{r})-2\sum_{r=1}^{k}|1-a_{r}|\lVert z_{r}\rVert>\frac{1}{2}\Delta^{i,j}(\sum_{r=1}^{k}z_{r}).$
and,
$\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})=$ $\frac{1}{\lVert w_{i}\rVert}|{\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}-{\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}|$
$\leq$ $\frac{1}{\lVert w_{i}\rVert}|{\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}z_{r}\rangle}-{\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}z_{r}\rangle}|+\sum_{r=1}^{k}2|1-a_{r}|\lVert z_{r}\rVert$
$\geq$ $\Delta^{i,j}(\sum_{r=1}^{k}z_{r})+2\sum_{r=1}^{k}|1-a_{r}|\lVert z_{r}\rVert\leq 2\Delta^{i,j}(\sum_{r=1}^{k}z_{r}).$
Therefore,
$\frac{1}{2}\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\leq\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\leq 2\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}).$
4.
Let $g_{1}\in\cup_{l=1}^{j}L^{i,l}(\sum_{r=1}^{k}z_{r})$. If $g_{1}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then we just showed that $g_{1}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. If $g_{1}\notin L^{i,j}(\sum_{r=1}^{k}z_{r})$ then for every $h\in\cup_{l\geq j}L^{i,l}L^{i,l}(\sum_{r=1}^{k}z_{r})$
${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}-{\langle U_{h}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}$ $\geq{\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}z_{r}\rangle}-\lVert w_{i}\rVert\lVert\sum_{r=1}^{k}|1-a_{r}|z_{r}\rVert$
$>\lVert w_{i}\rVert(\Delta^{i,j}(\sum_{r=1}^{k}z_{r})-\lVert\sum_{r=1}^{k}|1-a_{r}|z_{r}\rVert)$
$>0.$
But from 2.2 $|\cup_{l\geq j}L^{i,l}L^{i,l}(\sum_{r=1}^{k}z_{r})|\geq N-j+1$. So $g_{1}\in\cup_{l=1}^{j}L^{i,l}(\sum_{r=1}^{k}z_{r})$. Therefore,
$\cup_{l\leq j}L^{i,l}(\sum_{r=1}^{k}z_{r})\subset\cup_{l\leq j}L^{i,l}(\sum_{r=1}^{k}a_{r}z_{r}).$
The other inclusions are obtained similarly.
5. We prove the equality between complements:
$\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}=\left(L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\right)^{c}.$
First notice that by 2.3 we have that $L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$ and $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. From 2.6 part (2) we have that $L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. Take $g\in L^{i,j}(\sum_{r=1}^{k}z_{r}+e)$ and $h\in\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}$. Hence ${\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}+e\rangle}\neq{\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}+e\rangle}$.
There are two cases:
Case 1. $h\in L^{i,j}(\sum_{r=1}^{k}z_{r})\setminus L^{i,j}(\sum_{r=1}^{k}z_{r}+e)$. Thus ${\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}\rangle}={\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}\rangle}$. Therefore ${\langle U_{g}w_{i},e\rangle}\neq{\langle U_{h}w_{i},e\rangle}$. On the other hand $h\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$ since $L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. Hence ${\langle U_{g}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{h}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}$, which implies ${\langle U_{g}w_{i},\sum_{r=1}^{k}a_{r}z_{r}+e\rangle}\neq{\langle U_{h}w_{i},\sum_{r=1}^{k}a_{r}z_{r}+e\rangle}$. Thus $h\in\left(L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\right)^{c}$.
Case 2. $h\in G\setminus L^{i,j}(\sum_{r=1}^{k}z_{r})$. Thus ${\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}\rangle}\neq{\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}\rangle}$. In this case $|{\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}\rangle}-{\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}\rangle}|\geq\lVert w_{i}\rVert\Delta^{i,j}(\sum_{r=1}^{k}z_{r})$ and
$\left|{\langle U_{h}w_{i},\sum_{r=1}^{k}a_{r}z_{r}+e\rangle}-{\langle U_{g}w_{i},\sum_{r=1}^{k}a_{r}z_{r}+e\rangle}\right|\geq$
$\geq\left|{\langle U_{h}w_{i},\sum_{r=1}^{k}z_{r}\rangle}-{\langle U_{g}w_{i},\sum_{r=1}^{k}z_{r}\rangle}\right|$ $-\left|{\langle U_{h}w_{i},\sum_{r=1}^{k}(a_{r}-1)z_{r}+e\rangle}\right|$
$-\left|{\langle U_{g}w_{i},\sum_{r=1}^{k}(a_{r}-1)z_{r}+e\rangle}\right|\geq$
[…]
Hence again $h\in\left(L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\right)^{c}$. This proves that $\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}\subset\left(L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\right)^{c}$. The reverse inclusion is shown similarly, with $\Delta^{i,j}(\sum_{r=1}^{k}z_{r})$ replaced by $\Delta^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$ in (*). ∎
2.2.2 Positivity of the Lower Lipschitz Constant
Now we prove that the lower Lipschitz bound must be positive if the embedding map $\hat{\Phi}_{{\textbf{w}},S}$ is injective. We do so by contradiction.
The strategy is the following: Assume the lower Lipschitz constant is zero.
  • First we find a unit norm vector $z_{1}$ where the local lower Lipschitz constant vanishes.
  • Next we construct inductively a sequence of non-zero vectors $z_{2},z_{3},...,z_{k}$ so that the local lower Lipschitz constant vanishes in a convex set of the form $\{\sum_{r=1}^{k}a_{r}z_{r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ |a_{r}-1|<\delta\}$ for some $\delta>0$ small enough, and where sets $L^{i,j}$ remain constant. These steps are depicted in Figure 3 that suggests a "robotic arm" procedure (this name was suggested by the authors of [qaddura2024stable]).
  • For $k=d$ this construction defines a non-empty open set where the local lower Lipschitz constant vanishes and $L^{i,j}$ remain constants. This allows us to construct $u,v\neq 0$ so that $x=u+\sum_{r=1}^{d}z_{r}$ and $y=v+\sum_{r=1}^{d}z_{r}$ satisfy $x\not\sim y$ and yet $\Phi_{{\textbf{w}},S}(x)=\Phi_{{\textbf{w}},S}(y)$. This contradicts the injectivity hypothesis.
First, we show that if the lower bound is zero then it can be achieved locally. With a slightly abuse of notation we define
Lemma 2.8.
Fix ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$. If the lower Lipschitz constant of map $\Phi_{{\textbf{w}},S}$ is zero, then there exist sequences $(x_{n})_{n}$, $(y_{n})_{n}$ in ${\mathscr{V}}$ such that
$\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{d(x_{n},y_{n})^{2}}=0$
and, additionally, satisfy the following relations:
  1. 1.
    (convergence) They share a common limit $z_{1}$,
    $\lim_{n\rightarrow\infty}x_{n}=\lim_{n\rightarrow\infty}y_{n}=z_{1},$ (14)
    with $\lVert z_{1}\rVert=1$;
  2. 2.
    (boundedness) For all $k$:
    $\lVert x_{n}\rVert=1$ (15)
    $\lVert y_{n}\rVert\leq 1$ (16)
  3. 3.
    (alignment) For all $k$:
    $\lVert x_{n}-y_{n}\rVert=\min_{g\in G}\lVert x_{n}-U_{g}y_{n}\rVert$ (17)
    $\lVert x_{n}-z_{1}\rVert=\min_{g\in G}\lVert x_{n}-U_{g}z_{1}\rVert$ (18)
    $\lVert y_{n}-z_{1}\rVert=\min_{g\in G}\lVert y_{n}-U_{g}z_{1}\rVert$ (19)
Proof.
Because the lower Lipschitz bound of map $\Phi_{{\textbf{w}},S}$ is zero we have that
$\inf_{\begin{subarray}{c}x,y\in{\mathscr{V}}\\ x\nsim y\end{subarray}}\frac{\lVert\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)\rVert^{2}}{\operatorname{{\mathbf{d}}}([x],[y])^{2}}=0.$
Thus, we can find sequences $(x_{n})_{n},(y_{n})_{n}\in{\mathscr{V}}$ such that
$\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}=0$
Now, notice that for all $t>0$ we have $\Phi_{{\textbf{w}},S}(tx)=t\Phi_{{\textbf{w}},S}(x)$ and $\operatorname{{\mathbf{d}}}([tx],[ty])=t\operatorname{{\mathbf{d}}}([x],[y])$. So, for every $t>0$
$\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
$=\frac{\lVert\Phi_{{\textbf{w}},S}(tx_{n})-\Phi_{{\textbf{w}},S}(ty_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([tx_{n}],[ty_{n}])^{2}}.$
By setting $t=\frac{1}{\max(\lVert x_{n}\rVert,\lVert y_{n}\rVert)}$ we can always assume that both $x_{n}$ and $y_{n}$, lie in the unit ball, and what is more thanks to the symmetry of the formulas we can additionally assume that one of the sequences, say $x_{n}$, lies on unit sphere. In other words, $\lVert x_{n}\rVert=1$ and $\lVert y_{n}\rVert\leq 1$ for all $n\in\mathbb{N}$.
Because of this, we can find a convergent subsequence $(x_{n_{k}})_{k}$ of $(x_{n})_{n}$ with $x_{n_{k}}\to x_{\infty}$. Similarly, we can find a convergent subsequence $(y_{n_{k_{l}}})_{l}$ of $(y_{n_{k}})_{n}$ with $y_{n_{k_{l}}}\to y_{\infty}$. Clearly, $x_{n_{k_{l}}}\to x_{\infty}$. For easiness of notation, we denote the sequences $(x_{n_{k_{l}}})_{l}$ and $(y_{n_{k_{l}}})_{l}$ by $(x_{n})_{n}$ and $(y_{n})_{n}$, respectively.
Next, suppose that $x_{\infty}\nsim y_{\infty}$. Then,
$\frac{\lVert\Phi_{{\textbf{w}},S}(x_{\infty})-\Phi_{{\textbf{w}},S}(y_{\infty})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{\infty}],[y_{\infty}])^{2}}=\displaystyle\lim_{k\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}=0,$
and thus, $\Phi_{{\textbf{w}},S}(x_{\infty})=\Phi_{{\textbf{w}},S}(y_{\infty})$, which contradict the injectivity assumption. Hence, $x_{\infty}\sim y_{\infty}$.
Now, let us denote by $g_{\infty}$ a group element such that $x_{\infty}=U_{g_{\infty}}y_{\infty}$. Observe that $\lim_{n\to\infty}\lVert x_{n}-U_{g_{\infty}}y_{n}\rVert=0$. For each $n\in\mathbb{N}$ there exists at least one element $g_{n}\in G$, which achieves the Euclidean distance between $x_{n}$ and $U_{g_{\infty}}y_{n}$, i.e. satisfying $\operatorname{{\mathbf{d}}}([x_{n}],[U_{g_{\infty}}y_{n}])=\lVert x_{n}-U_{g_{k}g_{\infty}}y_{n}\rVert$. But $G$ is a finite group, meaning that, as $n$ goes to infinity, there must exist an element $g_{0}\in G$ for which $g_{n}=g_{0}$ for infinitely many $n$. Let $(n_{m})_{m}$ be the sequence of all such indices. We see that $\operatorname{{\mathbf{d}}}([x_{n_{m}}],[U_{g_{\infty}}y_{n_{m}}])=\lVert x_{n_{m}}-U_{g_{0}g_{\infty}}y_{n_{m}}\rVert$ for all $m\in\mathbb{N}$. Finally, for every $m\in\mathbb{N}$, let $g_{m}\in G$ be a group element that achieves the Euclidean distance between $x_{n_{m}}$ and $x_{\infty}$, that is
$\operatorname{{\mathbf{d}}}([x_{n_{m}}],[x_{\infty}])=\lVert U_{g_{m}}x_{n_{m}}-x_{\infty}\rVert.$
Denote $U_{g_{m}}x_{n_{m}}$ by $x_{n}$ and $U_{g_{m}g_{0}g_{\infty}}y_{n_{m}}$ by $y_{n}$. So far we obtained two sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ that satisfy (14-18). Now let $h_{n}\in G$ denote a group element so that $d(y_{n},z_{1})=\lVert y_{n}-U_{h_{n}}z_{1}\rVert$. Since $G$ is finite, pass to a subsequence (again indexed by $n$) so that $h_{n}=h_{0}$. Therefore $d(y_{n},z_{1})=\lVert y_{n}-U_{h_{0}}z_{1}\rVert\leq\lVert y_{k}-z_{1}\rVert$. But $\lim_{n\to\infty}y_{n}=z_{1}$. Thus $U_{h_{0}}z_{1}=z_{1}$. This shows (19) and the lemma is now proved. ∎
$z_{1}$$x$$y$$z$$x_{n}$$y_{n}$
Figure 2: Local analysis of sequences $x_{n}$ and $y_{n}$ converging to $z_{1}$.
In what follows, we will denote by $H(z)$ the stabilizer group of $z$; recall that
$H(z)=\{g\in G:U_{g}z=z\}.$
For a fixed vector $z$ we define the strictly positive number
$\rho_{0}(z)=\begin{cases}\min_{g\in G\setminus H(z)}\lVert z-U_{g}z\rVert,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ H(z)\neq G\\ \lVert z\rVert,\leavevmode\nobreak\ \text{if}\leavevmode\nobreak\ H(z)=G.\end{cases}$
Assume $N_{0}$ is large enough so that $d(x_{1,k},z_{1})<\frac{1}{8}\rho_{0}(z_{1})$ and $d(x_{1,k},y_{1,k})<\frac{1}{8}\rho_{0}(z_{1})$ for all $k>N_{0}$. Then
$\lVert y_{1,k}-z_{1}\rVert\leq\lVert y_{1,k}-x_{1,k}\rVert+\lVert x_{1,k}-z_{1}\rVert=d(y_{1,k},x_{1,k})+d(x_{1,k},z_{1})<\frac{\rho_{0}(z_{1})}{4}.$
Lemma 2.9.
Assume that $\lVert u\rVert,\lVert v\rVert<\frac{1}{4}\Delta_{0}(z_{1})$ and let $x=z_{1}+u$ and $y=z_{1}+v$. Then, the following properties hold:
  1. 1.
    $\operatorname{{\mathbf{d}}}([x],[z_{1}])=\lVert u\rVert$ and $\operatorname{{\mathbf{d}}}([y],[z_{1}])=\lVert v\rVert$,
  2. 2.
    $\operatorname{{\mathbf{d}}}([x],[y])=\min_{g\in H(z_{1})}\lVert u-U_{g}v\rVert=\min_{g\in H(z_{1})}\lVert U_{g}u-v\rVert$, and
  3. 3.
    the following are equivalent:
    1. (a)
      $\operatorname{{\mathbf{d}}}([x],[y])=\lVert u-v\rVert$,
    2. (b)
      $\lVert u-v\rVert\leq\lVert U_{g}u-v\rVert$, for all $g\in H(z_{1})$,
    3. (c)
      ${\langle u,v\rangle}\geq{\langle U_{g}u,v\rangle}$, for all $g\in H(z_{1})$.
Proof.
  1. 1.
    If $u=0$ then the claim follows. If $u\not=0$, then $\operatorname{{\mathbf{d}}}([x],[z_{]}1)=\min_{g\in G}\lVert x-U_{g}z_{1}\rVert=\min_{g\in G}\lVert z_{1}-U_{g}z_{1}+u\rVert\leq\lVert u\rVert$. From the other hand, suppose that minimum is achieved for a permutation $g\in G$. If $g\in H(z_{1})$, then $\operatorname{{\mathbf{d}}}([x],[z_{1}])=\lVert u\rVert$. If $g\notin H(z_{1})$, then $\operatorname{{\mathbf{d}}}([x],[z_{1}])>\lVert u\rVert\leq\operatorname{{\mathbf{d}}}([x],[z_{1}])$, which is a contradiction.
  2. 2.
    Obviously $\operatorname{{\mathbf{d}}}([x],[z_{1}])\leq\min_{g\in H(z_{1})}\lVert U_{g}u-v\rVert$. On the other hand, for $g\in G\setminus K$ and $h\in G$,
    $\lVert U_{g}x-y\rVert$ $=\lVert U_{g}z_{1}-z_{1}+U_{g}u-v\rVert$
    $\geq\lVert U_{g}z_{1}-z_{1}\rVert-\lVert u\rVert-\lVert v\rVert$
    $\geq\rho_{0}(z_{1})-2\lVert u\rVert-2\lVert v\rVert+\lVert U_{h}u-v\rVert$
    $\geq\operatorname{{\mathbf{d}}}([x],[y]).$
  3. 3.
    • $(a)\Rightarrow(b)$. If $\operatorname{{\mathbf{d}}}([x],[y])=\lVert u-v\rVert$, then $\lVert u-v\rVert\leq\lVert U_{g}x-y\rVert=\lVert U_{g}z_{1}-z_{1}+U_{g}u-v\rVert$, $\forall g\in G$. For $g\in H(z_{1})$ this reduces to (b)
    • $(b)\Rightarrow(a)$. Assume that $\forall g\in H(z_{1})$, $\lVert u-v\rVert\leq\lVert U_{g}u-v\rVert$ Then $\lVert u-v\rVert=\lVert x-y\rVert\leq\lVert U_{g}u-v\rVert=\lVert U_{g}x-y\rVert$ For, $g\in G\setminus H(z_{1})$ $\lVert U_{g}x-y\rVert=\lVert U_{g}z_{1}-z_{1}+U_{g}u-v\rVert\geq\lVert U_{g}z_{1}-z_{1}\rVert-\lVert u\rVert-\lVert v\rVert\geq\rho_{0}(z_{1})-2\lVert u\rVert-2\lVert v\rVert+\lVert u-v\rVert\geq\lVert u-v\rVert=\lVert x-y\rVert$ Thus, $\operatorname{{\mathbf{d}}}([x],[y])=\lVert x-y\rVert=\lVert u-v\rVert$
    • $(b)\Leftrightarrow(c)$ is immediate from definition of inner product
Remark 2.10.
Applying Lemma 2.9 to two sequences $(x_{k})_{k}$ and $(y_{k})_{k}$ that satisfy (14-17) in 2.8, it follows that $d(x_{k},z_{1})=\lVert x_{k}-z_{1}\rVert$ and $d(y_{k},z_{1})=\lVert y_{k}-z_{1}\rVert$ for $k$ large enough. Hence alignment must occur from some rank on.
Lemma 2.11.
For fixed $i\in[p]$, $j\in S_{i}$ and two sequences $(x_{n})_{n}$, $(y_{n})_{n}$ produced by 2.8, we denote by $g_{1,n,i,j}$ the group elements that achieves $\Phi_{i,j}(x_{n})$ and by $g_{2,n,i,j}$ the group element that achieves the $\Phi_{i,j}^{j}(y_{n})$. That is $\Phi_{i,j}(x_{n})={\langle U_{g_{1,n,i,j}}w_{i},x_{n}\rangle}$ and $\Phi_{i,j}(y_{n})={\langle U_{g_{2,n,i,j}}w_{i},y_{n}\rangle}$.
We can find a sequence of natural numbers $(n_{r})_{r}$, such that, $g_{1,n_{r},i,j}=g_{1,i,j}$ and $g_{2,n_{r},i,j}=g_{2,i,j}$ $\forall r\in\mathbb{N},\leavevmode\nobreak\ i\in[p],\leavevmode\nobreak\ j\in S_{i}$.
Proof.
For $i=1,\leavevmode\nobreak\ j=1$ there is a subsequence $(x_{n_{m}})_{m}$ such that $g_{1,1,1,n_{m}}=g_{1,1,1}$ for every $m\in\mathbb{N}$. Similarly, for $i=1,\leavevmode\nobreak\ j=2$ we can find a subsequence of $(x_{n_{m}})_{m}$, lets call it $(x_{n_{l}})_{l}$, such that $g_{1,1,2,n_{l}}=g_{1,1,2},\leavevmode\nobreak\ \forall l\in\mathbb{N}$. So by induction after $\sum_{i\in[p]}m_{i}=m$ steps we construct a subsequence of $(x_{n})_{n}$ lets call it $(x_{n_{m}})_{m}$ such that $g_{1,i,j,n_{m}}=g_{1,i,j}$ for every $i\in[p],j\in S_{i}$. Starting from sequence $(y_{1,n_{m}})_{m}$ we repeat the same procedure concluding in a subsequence $(y_{1,n_{r}})_{r}$ such that $g_{2,i,j,n_{r}}=g_{2,i,j}$ for every $r\in\mathbb{N},i\in[p],j\in S_{i}$ Notice that sequences $(x_{n_{r}})_{r}$ and $(y_{n_{r}})_{r}$ that from now on we will call them $(x_{n})_{n}$ and $(y_{n})_{n}$ for easiness of notation, satisfy the assumptions of lemma. ∎
For sequences $(x_{n})_{n}$, $(y_{n})_{n}$ and $z_{1}$ defined before, let $u_{n}=x_{n}-z_{1}$ and $v_{n}=y_{n}-z_{1}$. Notice that
$\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}=\sum_{i=1}^{p}\sum_{j\in S_{i}}|$ ${\langle U_{g_{1,i,j}}w_{i},x_{n}\rangle}-{\langle U_{g_{2,i,j}}w_{i},y_{n}\rangle}|^{2}$
$=\sum_{i=1}^{p}\sum_{j\in S_{i}}|$ ${\langle U_{g_{1,i,j}}w_{i}-U_{g_{2,i,j}}w_{i},z_{1}\rangle}$
$+$ ${\langle w_{i},U_{g_{1,i,j}^{-1}}u_{n}-U_{g_{2,i,j}^{-1}}v_{n}\rangle}|^{2}.$
This sequence converge to 00, as $k\to\infty$ while also $u_{n},v_{n}\to 0$. So we conclude that for each $i\in[p]$ and $j\in S_{i}$, ${\langle U_{g_{1,i,j}}w_{i}-U_{g_{2,i,j}}w_{i},z_{1}\rangle}=0$. So
$\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}=\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle w_{i},U_{g_{1,i,j}^{-1}}u_{n}-U_{g_{2,i,j}^{-1}}v_{n}\rangle}|^{2}.$
Thus we have
$\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle w_{i},U_{g_{1,i,j}^{-1}}u_{n}-U_{g_{2,i,j}^{-1}}v_{n}\rangle}|^{2}}{\lVert u_{n}-v_{n}\rVert^{2}}=0$ (20)
where $\lVert u_{n}\rVert,\lVert v_{n}\rVert\to 0$, so for large enough $n$ we have that $\lVert u_{n}\rVert,\lVert v_{n}\rVert\leq\frac{1}{4}\rho_{0}(z_{1})$. Recall that from 2.9, we conclude that exists $N_{0}\in\mathbb{N}$, such that $\lVert u_{n}-v_{n}\rVert\leq\lVert U_{g}u_{n}-v_{n}\rVert$ for all $g\in H(z_{1})$ and $k\geq N_{0}$.
Lemma 2.12.
Fix $p\in\mathbb{N}$, ${\textbf{w}}\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$. Let $\Delta:{\mathscr{V}}\to\mathbb{R}$, where $\Delta(x)=\min_{(i,j)\in[p]\times[N]}\Delta^{i,j}(x)$, where the map $\Delta^{i,j}$ is defined in (5). Fix nonzero vectors $z_{1},\dots,z_{k}\in{\mathscr{V}}$, such that
$\lVert z_{1}\rVert=1,\leavevmode\nobreak\ {\langle z_{i},z_{j}\rangle}=0,\leavevmode\nobreak\ \forall i,j\in[k],\leavevmode\nobreak\ i\not=j$
and
$\lVert z_{l+1}\rVert\leq\min\left(\frac{1}{4}\Delta(\sum_{r=1}^{l}z_{r}),\frac{1}{4}\lVert z_{l}\rVert\right),\leavevmode\nobreak\ \forall l\in[k-1].$
Assume that the local lower Lipschitz constant of $\Phi_{{\textbf{w}},S}$ vanishes at $z_{1}+z_{2}+\cdots+z_{k}$.
  1. 1.
    The local lower Lipschitz constant vanishes on the non-empty convex box $\{\sum_{r=1}^{k}a_{r}z_{r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ |a_{r}-1|<\frac{1}{16k}\Delta(\sum_{l=1}^{k}z_{l})\}$ centered at $z_{1}+z_{2}+\cdots+z_{k}$.
  2. 2.
    Assume $\hat{\Phi}_{{\textbf{w}},S}$ is injective. If $k<d$ then there exists a nonzero vector $z_{k+1}$ such that:
    (i) ${\langle z_{k+1},z_{j}\rangle}=0,\leavevmode\nobreak\ \forall j\in[k]$;
    (ii) $\lVert z_{k+1}\rVert\leq\min\left(\frac{1}{4}\Delta(\sum_{r=1}^{k}z_{r}),\frac{1}{4}\lVert z_{k}\rVert\right)$; and
    (iii) The local lower Lipschitz constant vanishes at $z_{1}+z_{2}+\cdots+z_{k+1}$, i.e. there are sequences of vectors $(x_{n})_{n},(y_{n})_{n}$ such that
    $\lim_{n\to\infty}x_{n}=\lim_{n\to\infty}y_{n}=\sum_{r=1}^{k+1}z_{r}$
    and
    $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}=0.$
Proof.
  1. 1.
    Let $(x_{n})_{n},(y_{n})_{n}$ be sequences in ${\mathscr{V}}$ such that
    $\lim_{n\to\infty}x_{n}=\lim_{n\to\infty}y_{n}=\sum_{r=1}^{k}z_{r}$
    and
    $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}=0.$
    Claim: For any $a_{1},\dots,a_{k}\in\left(1-\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r}),1+\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})\right)$ the sequences
    $\tilde{x}_{n}=x_{n}+\sum_{r=1}^{k}(a_{r}-1)z_{r}$
    and
    $\tilde{y}_{n}=y_{n}+\sum_{r=1}^{k}(a_{r}-1)z_{r}$
    also achieve a zero lower Lipschitz constant, i.e.
    $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([\tilde{x}_{n}],[\tilde{y}_{n}])^{2}}=0.$
    First we denote by $u_{n}$ and $v_{n}$ the difference sequences $x_{n}$ and $y_{n}$ to their common limit $\sum_{r=1}^{k}z_{r}$,
    $u_{n}=x_{n}-\sum_{r=1}^{k}z_{r}=\tilde{x}_{n}-\sum_{r=1}^{k}a_{r}z_{r}$
    and
    $v_{n}=y_{n}-\sum_{r=1}^{k}z_{r}=\tilde{y}_{n}-\sum_{r=1}^{k}a_{r}z_{r}.$
    Sequences $(u_{n})_{n}$ and $(v_{n})_{n}$ converge to zero. Therefore there exists $M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$
    1. (a)
      $\lVert u_{n}\rVert=\lVert x_{n}-\sum_{r=1}^{k}z_{r}\rVert<\frac{1}{16}\Delta(\sum_{r=1}^{k}z_{r})$
    2. (b)
      $\lVert u_{n}\rVert=\lVert\tilde{x}_{n}-\sum_{r=1}^{k}a_{r}z_{r}\rVert<\frac{1}{16}\Delta(\sum_{r=1}^{k}a_{r}z_{r})$
    3. (c)
      $\lVert v_{n}\rVert=\lVert y_{n}-\sum_{r=1}^{k}z_{r}\rVert<\frac{1}{16}\Delta(\sum_{r=1}^{k}z_{r})$
    4. (d)
      $\lVert v_{n}\rVert=\lVert\tilde{y}_{n}-\sum_{r=1}^{k}a_{r}z_{r}\rVert<\frac{1}{16}\Delta(\sum_{r=1}^{k}a_{r}z_{r})$.
    Thus from part (3) of 2.6, 2.3 and part (2) of 2.6 we have that for any $n\geq M_{0}$ and $(i,j)\in S$
    $L^{i,j}(\tilde{x}_{n})=L^{i,j}(x_{n})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$
    and
    $L^{i,j}(\tilde{y}_{n})=L^{i,j}(y_{n})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}).$
    Therefore,
    $0=$ $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}=$
    $=$ $\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle w_{i},U_{g_{1,i,j}^{-1}}u_{n}-U_{g_{2,i,j}^{-1}}v_{n}|^{2}\rangle}}{\lVert u_{n}-v_{n}^{2}\rVert}=$
    $=$ $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([\tilde{x}_{n}],[\tilde{y}_{n}])^{2}},$
    where
    $g_{1,i,j}\in L^{i,j}(x_{n})=L^{i,j}(\tilde{x}_{n})$
    and
    $g_{2,i,j}\in L^{i,j}(y_{n})=L^{i,j}(\tilde{y}_{n}).$
    This proves the lower Lipschitz constant of $\Phi_{{\textbf{w}},S}$ vanishes at $\sum_{r=1}^{k}a_{r}z_{r}$.
  2. 2.
    Let two sequences $(x_{n})_{n},(y_{n})_{n}$ that both converge to $\sum_{r=1}^{k}z_{r}$, and achieve lower Lipschitz bound zero for map $\Phi_{{\textbf{w}},S}$. We align sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ to satisfy the properties of 2.8. We denote by $a_{n}=P_{E_{k}}x_{n}$ and $b_{n}=P_{E_{k}}y_{n}$ the orthogonal projections of the sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ respectively, on the linear subspace $E_{k}=\operatorname{span}\{z_{1},\dots,z_{k}\}^{\bot}$.
    Claim 1: First we will show that $\exists M_{0}$ such that $\forall n\geq M_{0}$, $a_{n}\neq 0$ or $b_{n}\neq 0$. Assuming otherwise, there are two sequences of vectors $x_{n}=\sum_{r=1}^{k}c_{r,n}z_{r}$ and $y_{n}=\sum_{r=1}^{k}d_{r,n}z_{r}$, where $\lim_{n\to\infty}c_{r,n}=\lim_{n\to\infty}d_{r,n}=1,\leavevmode\nobreak\ \forall r\in[k]$ that achieve lower Lipschitz bound zero. Recall that from part (2) of 2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
    $L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{k}d_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{k}z_{r}).$
    Then, for $g_{i,j}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$,
    00 0 $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
    $=\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},x_{n}-y_{n}\rangle}|^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
    $\geq\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},\sum_{r=1}^{k}(c_{r,n}-d_{r,n})z_{r}\rangle}|^{2}}{\lVert\sum_{r=1}^{k}(c_{r,n}-d_{r,n})z_{r}\rVert^{2}}$
    $=\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},\tilde{z}\rangle}|^{2},$
    where
    $\tilde{z}=\lim_{m\to\infty}\frac{\sum_{r=1}^{k}(c_{r,n_{m}}-d_{r,{n_{m}}})z_{r}}{\lVert\sum_{r=1}^{k}(c_{r,n_{m}}-d_{r,{n_{m}}})z_{r}\rVert}$
    is a unit vector obtained as the limit of a convergent subsequence of the sequence of unit vectors $\frac{\sum_{r=1}^{k}(c_{r,n}-d_{r,{n}})z_{r}}{\lVert\sum_{r=1}^{k}(c_{r,n}-d_{r,n})z_{r}\rVert}$. Since the group $G$ is finite, we can find a positive number $\epsilon>0$ such that $\epsilon\lVert\tilde{z}\rVert<\frac{1}{4}\Delta(\sum_{r=1}^{k}z_{r})$ and $\sum_{r=1}^{k}z_{r}\nsim\sum_{r=1}^{k}z_{r}+\epsilon\tilde{z}$. In this case
    $\Phi_{{\textbf{w}},S}(\sum_{r=1}^{k}z_{r})=\Phi_{{\textbf{w}},S}(\sum_{r=1}^{k}z_{r}+\epsilon\tilde{z})$
    which contradict the injectivity property. This establishes Claim 1.
    Now we can assume for all $n\geq M_{0}$, $a_{n}=P_{E_{k}}x_{n}\neq 0$ or $b_{n}=P_{E_{k}}y_{n}\neq 0$. If need be, pass to a subsequence and/or switch the definitions of $x_{n}$ and $y_{n}$, so that $\lVert b_{n}\rVert\geq\lVert a_{n}\rVert$ for all $n$. In doing so we no longer claim the normalization (15). Nevertheless, both $\lVert x_{n}\rVert,\lVert y_{n}\rVert\leq 1$.
    Let $c_{r,n},d_{r,n}$ be the unique coefficients determined by $x_{n}=\sum_{r=1}^{k}c_{r,n}z_{r}+a_{n}$, $y_{n}=\sum_{r=1}^{k}d_{r,n}z_{r}+b_{n}$. Note $\lim_{n\rightarrow\infty}c_{r,n}=\lim_{n\rightarrow\infty}d_{r,n}=1$.
    Let $e_{n}=\sum_{r=1}^{k}(d_{r,n}-c_{r,n})z_{r}+b_{n}$ and
    $s_{n}=\frac{\min\left(\lVert z_{k}\rVert,\Delta(\sum_{r=1}^{k}z_{r}),\rho_{0}(\sum_{r=1}^{k}z_{r})\right)}{16\lVert e_{n}\rVert}.$
    Note $\lVert e_{n}\rVert\geq\lVert b_{n}\rVert\geq\lVert a_{n}\rVert$ for all $n$.
    Claim 2: Sequences $\tilde{x}_{n}=\sum_{r=1}^{k}z_{r}+s_{n}a_{n}$ and $\tilde{y}_{n}=\sum_{r=1}^{k}z_{r}+s_{n}e_{n}$ achieve also the lower Lipschitz constant zero at $\sum_{r=1}^{k}z_{k}$.
    Note that $\max(\lVert s_{n}a_{n}\rVert,\lVert s_{n}e_{n}\rVert)\leq\frac{1}{16}$. Pass to subsequences of $(a_{n})_{n}$ and $(e_{n})_{n}$ so that both $\lim_{n\to\infty}s_{n}a_{n}$ and $\lim_{n\to\infty}s_{n}e_{n}$ converge. Let $\alpha=\lim_{n\to\infty}s_{n}a_{n}$ and $\delta=\lim_{n\to\infty}s_{n}e_{n}$. Notice $\delta\neq 0$.
    The limits
    $\lim_{n\to\infty}c_{r,n}=\lim_{n\to\infty}d_{r,n}=1,\leavevmode\nobreak\ \forall r\in[k]\leavevmode\nobreak\ \text{and }\lim_{n\to\infty}a_{n}=\lim_{n\to\infty}e_{n}=0$
    imply that $\exists m_{0}\in\mathbb{N}$ such that $\forall n\geq m_{0}$, and $\forall r\in[k]$
    1. (a)
      $|1-c_{r,n}|<\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})$
    2. (b)
      $|1-d_{r,n}|<\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})$
    3. (c)
      $|c_{r,n}-d_{r,n}|<\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})$
    4. (d)
      $\lVert a_{n}\rVert<\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})$
    5. (e)
      $\lVert e_{n}\rVert<\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})$
    From 2.6 part (1),
    $\Delta(\sum_{r=1}^{k}c_{r,n}z_{r})\geq\frac{1}{4}\Delta(\sum_{r=1}^{k}z_{r}).$
    Also
    $\max(\lVert a_{n}\rVert,\lVert s_{n}a_{n}\rVert)<\frac{1}{16}\Delta(\sum_{r=1}^{k}z_{r})\leq\frac{1}{4}\Delta(\sum_{r=1}^{k}c_{r,n}z_{r})$
    and
    $\max(\lVert e_{n}\rVert,\lVert s_{n}e_{n}\rVert)<\frac{1}{16}\Delta(\sum_{r=1}^{k}z_{r})\leq\frac{1}{4}\Delta(\sum_{r=1}^{k}d_{r,n}z_{r})$
    So, for any $(i,j)\in S$
    $L^{i,j}(x_{n})$ $=L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r}+a_{n})=L^{i,j}(\sum_{r=1}^{k}z_{r}+a_{n})$
    $=L^{i,j}(\sum_{r=1}^{k}z_{r}+s_{n}a_{n})=L^{i,j}(\tilde{x}_{n})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r}).$
    Where the second equality comes from 2.6 part 3, third equality from 2.4 the fifth inclusion from 2.3, and the last equality from 2.6 part 2.
    Similarly,
    $L^{i,j}(y_{n})$ $=L^{i,j}(\sum_{r=1}^{k}d_{r,n}z_{r}+b_{n})=L^{i,j}(\sum_{r=1}^{k}(1+d_{r,n}-c_{r,n})z_{r}+b_{n})$
    $=L^{i,j}(\sum_{r=1}^{k}z_{r}+e_{n})=L^{i,j}(\sum_{r=1}^{k}z_{r}+s_{n}e_{n})=L^{i,j}(\tilde{y}_{n})$
    $\subset L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r}).$
    Therefore,
    $0=$ $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}=$
    $=$ $\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}{\langle w_{i},U_{g_{1,i,j}^{-1}}a_{n}-U_{g_{2,i,j}^{-1}}e_{n}\rangle}^{2}}{\lVert a_{n}-e_{n}\rVert^{2}}$
    $=$ $\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}{\langle w_{i},U_{g_{1,i,j}^{-1}}s_{n}a_{n}-U_{g_{2,i,j}^{-1}}s_{n}e_{n}^{2}\rangle}}{\lVert s_{n}a_{n}-s_{n}e_{n}\rVert^{2}}$
    $=$ $\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([\tilde{x}_{n}],[\tilde{y}_{n}])^{2}}.$
    where
    ${g_{1,i,j}}\in L^{i,j}(x_{n})\text{ and }{g_{2,i,j}}\in L^{i,j}(y_{n})$
    are chosen independent of $n$ by possibly passing to subsequences since $G$ is finite. So,
    $\Phi_{{\textbf{w}},S}(\sum_{r=1}^{k}z_{r}+\alpha)-\Phi_{{\textbf{w}},S}(\sum_{r=1}^{k}z_{r}+\delta)=0.$
    Since $\hat{\Phi}_{{\textbf{w}},S}$ is injective,
    $\sum_{r=1}^{k}z_{r}+\alpha\sim\sum_{r=1}^{k}z_{r}+\delta$
    Let $g_{1}\in G$ denote a group element that achieves this equivalence, i.e.
    $\sum_{r=1}^{k}z_{r}+\alpha=U_{g_{1}}(\sum_{r=1}^{k}z_{r}+\delta)$
    Note that $g_{1}\in H(\sum_{r=1}^{k}z_{r})$ because otherwise
    $0=\lVert\sum_{r=1}^{k}z_{r}+\alpha-U_{g_{1}}(\sum_{r=1}^{k}z_{r})+U_{g_{1}}\delta)\rVert=\lVert\sum_{r=1}^{k}z_{r}+\alpha-U_{g_{1}}(\sum_{r=1}^{k}z_{r})+U_{g_{1}}\delta)\rVert$
    $\geq\lVert\sum_{r=1}^{k}z_{r}-U_{g_{1}}(\sum_{r=1}^{k}z_{r})\rVert-\lVert\alpha-U_{g_{1}}\delta\rVert\geq\rho_{0}(\sum_{r=1}^{k}z_{r})-\lVert\alpha\rVert-\lVert\delta\rVert>0$
    The last inequality comes from the fact that $\lVert\alpha\rVert<\frac{1}{4}\rho_{0}(\sum_{r=1}^{k}z_{r})$, and $\lVert\delta\rVert<\frac{1}{4}\rho_{0}(\sum_{r=1}^{k}z_{r})$.
    Additionally, $\alpha=U_{g_{1}}\delta$ because
    $0=\lVert\sum_{r=1}^{k}z_{r}+a-U_{g_{1}}(\sum_{r=1}^{k}z_{r})+U_{g_{1}}\delta)\rVert=\lVert a-U_{g_{1}}\delta\rVert.$
    Claim 3: The two vectors $\alpha$ and $\delta$ are equal, $\alpha=\delta$.
    We prove this claim by contradiction. Assume that $\alpha\neq\delta$. From 2.9, $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$
    $\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert.$
    Therefore,
    $00<\lVert a-\delta\rVert=\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert=$ 0 .
    We conclude that $\alpha=\delta\neq 0$.
    Set $z_{k+1}=\alpha=\delta$. Together with sequences $\tilde{x}_{n}$ and $\tilde{y}_{n}$, they satisfy the assertions of part 2 of this Lemma.
Remark 2.13.
Our construction produces $z_{k+1}$ that has norm equal to
$\frac{1}{16}\min\left(\lVert z_{k}\rVert,\Delta(\sum_{r=1}^{k}z_{r}),\rho_{0}(\sum_{r=1}^{k}z_{r})\right)$.
Now we can complete the proof of 1.4.
$z_{1}$$z_{2}$
$z_{1}$$z_{2}$$z_{3}$
Figure 3: The robotic arm method.
Proof.
Starting from vector $z_{1}$ and the sequences $(x_{n})_{n}$, $(y_{n})_{n}$ observed in 2.1 after $d-1$ steps of algorithmic construction of part (2) of 2.12 we get $d$ non-zero vectors $\{z_{1},\dots,z_{d}\}$ and a pair of sequences $(\tilde{x}_{n})_{n}$, $(\tilde{y}_{n})_{n}$ such that
(i) ${\langle z_{i},z_{j}\rangle}=0,\leavevmode\nobreak\ \forall i,j\in[d],i\neq j$;
(ii) $\lVert z_{k+1}\rVert\leq\min(\frac{1}{4}\Delta(\sum_{r=1}^{k}z_{r}),\frac{1}{4}\lVert z_{k}\rVert),\leavevmode\nobreak\ \forall k\in[d-1];$ and
(iii) $\lim_{n\to\infty}\tilde{x}_{n}=\lim_{n\to\infty}\tilde{y}_{n}=\sum_{r=1}^{d}z_{r}$ and
$\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([\tilde{x}_{n}],[\tilde{y}_{n}])^{2}}=0.$
Let $\tilde{x}_{n}=\sum_{r=1}^{d}l_{r,n}z_{r}$ and $\tilde{y}_{n}=\sum_{r=1}^{d}t_{r,n}z_{r}$. Notice that $\lim_{n\to\infty}l_{r,n}=\lim_{n\to\infty}t_{r,n}=1,\leavevmode\nobreak\ \forall r\in[d]$.
Recall that from part (2) of 2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
$L^{i,j}(\sum_{r=1}^{d}l_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{d}t_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{d}z_{r}).$
Then, for $g_{i,j}\in L^{i,j}(\sum_{r=1}^{d}z_{r})$,
00 0 $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
$=\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},\tilde{x}_{n}-\tilde{y}_{n}\rangle}|^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
$\geq\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},\sum_{r=1}^{k}(l_{r,n}-t_{r,n})z_{r}\rangle}|^{2}}{\lVert\sum_{r=1}^{d}(l_{r,n}-t_{r,n})z_{r}\rVert^{2}}$
$=\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},\tilde{z}\rangle}|^{2},$
where
$\tilde{z}=\lim_{m\to\infty}\frac{\sum_{r=1}^{d}(l_{r,n_{m}}-d_{r,{n_{m}}})z_{r}}{\lVert\sum_{r=1}^{d}(l_{r,n_{m}}-t_{r,{n_{m}}})z_{r}\rVert}$
is a unit vector obtained as the limit of a convergent subsequence of the sequence of unit vectors $\frac{\sum_{r=1}^{d}(l_{r,n}-t_{r,{n}})z_{r}}{\lVert\sum_{r=1}^{d}(l_{r,n}-t_{r,n})z_{r}\rVert}$. Since the group $G$ is finite, we can find a positive number $\epsilon>0$ such that $\epsilon\lVert\tilde{z}\rVert<\frac{1}{4}\Delta(\sum_{r=1}^{d}z_{r})$ and $\sum_{r=1}^{d}z_{r}\nsim\sum_{r=1}^{d}z_{r}+\epsilon\tilde{z}$. In this case
$\Phi_{{\textbf{w}},S}(\sum_{r=1}^{d}z_{r})=\Phi_{{\textbf{w}},S}(\sum_{r=1}^{d}z_{r}+\epsilon\tilde{z})$
which contradict the injectivity property. 1.4 is now proved.
3 Dimension reduction using linear maps
In previous section we considered an embedding $\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\rightarrow\mathbb{R}^{m}$ that is injective on the quotient space $\hat{{\mathscr{V}}}$. The dimension $m$ of the space $\mathbb{R}^{m}$ may be very large. In this section we show that the nonlinear map $\hat{\Phi}_{{\textbf{w}},S}$ can be further linearly processed into a smaller dimensional space while preserving injectivity and bi-Lipschitz properties.
The idea of using dimension reduction linear maps goes back many years. The famous Johnson-Lindenstrauss Lemma [johnson1986extensions] provides a nearly isometric projection for finite metric spaces. More recently, [Cahill19] shows that a dimension reduction linear map preserves Lipschitz properties for translation invariant polynomial embeddings.
The first result that combines sorted co-orbits with linear maps was shown in [balan2022permutation] in the case of the group of permutations $G=S_{n}$ acting by left multiplication on $n\times d$ real matrices, ${\mathscr{V}}=\mathbb{R}^{n\times d}$. In that particular case, the dimension of the intermediary co-orbit space was $n(1+(d-1)n!)$ while the final target space was shown to be $2nd=2\,dim({\mathscr{V}})$. The authors of [dym2022low] show that intermediate embedding can be realized in dimension $(2nd+1)n$ instead of $n(1+(d-1)n!)$ with the final target space of dimension $12nd+$ 1. A more careful analysis of such an embedding by Matthias Wellershoff [Matthias2023a] proved that $12nd+$ 1 can be replaced by $2nd-d$. In the Ph.D. thesis [tsoukanis2024g], one of the authors showed that the decrease of dimension from $2nd$ to $2nd-d$ represents the dimension of a certain joint eigenspace of this representation.
The exact statement of our dimension reduction result is included in 1.6.
To prove 1.6 we need first to construct a few objects. Let $D:{\mathscr{V}}\times{\mathscr{V}}\to\mathbb{R}^{m}$ be the nonlinear map $D(x,y)=\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)$. Its range $E$ is defined by $E=\operatorname{Ran}(D):=\{\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y):x,y\in{\mathscr{V}}\}=\operatorname{Ran}(\Phi_{{\textbf{w}},S})-\operatorname{Ran}(\Phi_{{\textbf{w}},S})$.
Fix $g_{1},\dots,g_{N}$ an enumeration of the group elements. Define $\lambda_{i,j}(x):{\mathscr{V}}\to\mathbb{R}$ by $\lambda_{i,j}(x)={\langle U_{g_{i}}w_{j},x\rangle}$. Notice that $\lambda_{i,j}$ is a linear map (unlike $\Phi_{i,j}$) and also that
$\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)$ $=[\lambda_{1,\nu_{1}(1)}(x)-\lambda_{1,\nu_{p+1}(1)}(y),\dots,\lambda_{1,\nu_{1}(m_{1})}(x)-\lambda_{1,\nu_{p+1}(m_{1})}(y),$
$\dots,\lambda_{p,\nu_{p}(1)}(x)-\lambda_{p,\nu_{2p}(1)}(y),\dots,\lambda_{p,\nu_{p}(m_{p})}(x)-\lambda_{p,\nu_{2p}(m_{p})}(y)]$
for some permutations $\nu_{1},...,\nu_{2p}\in S_{N}$ that may depend on $x$ and $y$. Let $m_{j}=|S_{j}|=|\{i\in[N]\leavevmode\nobreak\ ,\leavevmode\nobreak\ (i,j)\in S\}|$ so that $m_{1}+\cdots+m_{p}=m$.
Now, fix permutations $\pi_{1},\ldots,\pi_{2p}\in S_{N}$ and let $L_{\pi_{1},\dots,\pi_{2p}}:{\mathscr{V}}\times{\mathscr{V}}\to\mathbb{R}^{m}$ denote the linear map
$L_{\pi_{1},\dots,\pi_{2p}}(x,y)=$ $[\lambda_{1,\pi_{1}(1)}(x)-\lambda_{1,\pi_{p+1}(1)}(y),\dots,\lambda_{1,\pi_{1}(m_{1})}(x)-\lambda_{1,\pi_{p+1}(m_{1})}(y),$
$\dots,\lambda_{p,\pi_{p}(1)}(x)-\lambda_{p,\pi_{2p}(1)}(y),\dots,\lambda_{p,\pi_{p}(m_{p})}(x)-\lambda_{p,\pi_{2p}(m_{p})}(y)]$
Define
$F=\cup_{\pi_{1},\dots,\pi_{2p}\in S_{N}}\operatorname{Ran}(L_{\pi_{1},\dots\pi_{2p}}).$
Notice that $F$ is a finite union of linear subspaces and that $E\subset F$. For fixed $\pi_{1},\dots\pi_{2p}$ the map $(x,y)\mapsto L_{\pi_{1},\dots,\pi_{2p}}(x,y)$ is linear in $(x,y)$, $L_{\pi_{1},\dots,\pi_{2p}}(v,v)=0$ for all $v\in V_{G}$, and from the rank-nullity theorem we have
$\dim(\operatorname{Ran}(L_{\pi_{1},\dots\pi_{2p}}))\leq 2d-d_{G}.$
Lemma 3.1.
Assume $r,s,m$ are non-negative integers so that $r+s\leq m$. For any finite collection $\{F_{a}\,:\,a\in[T]\}$ of $T$ linear subspaces of $\mathbb{R}^{m}$ of dimension at most $s$, a generic $r$-dimensional linear subspace $K$ of $\mathbb{R}^{m}$ , satisfies $K\cap F_{a}=\{0\},\ \forall a\in[T]$. Here generic means open and dense with respect to Zarisky topology.
Proof.
Let $\{v_{1},\dots v_{r}\}$ be a spanning set for $K$, and $\{w_{1,a}\dots w_{M-r,a}\}$ be a linearly independent set of vectors such that $F_{a}\subset\operatorname{span}\{w_{1,a},\dots,w_{m-r,a}\}$. Then, $\operatorname{span}\{v_{1},\dots,v_{r}\}\cap\operatorname{span}\{w_{1,a},\dots,w_{m-r,a}\}=\{0\}$ if, and only if, the set $\{v_{1},\dots,v_{r},w_{1,a},\dots,w_{m-r,a}\}$ is linearly independent. Define $R_{a}(v_{1},\dots v_{r})=\det[v_{1}|\dots v_{r}|w_{1,a}|\dots w_{m-r,a}]$, and note that $R_{a}(v_{1},\dots v_{r})$ is a polynomial in $rm$ variables $v_{1}(1),\dots,v_{1}(M),\dots v_{r}(1),\dots,v_{r}(m)$. Hence,
$K\cap F_{a}=\{0\},\quad\forall a\in[N]$ $\iff R_{a}(v_{1},\dots v_{r})\not=0,\quad\forall\ a\in[N]$
$\iff\prod_{a=1}^{N}R_{a}(v_{1},\dots,v_{r})\not=0.$
We conclude that
$\mathbb{U}=\left\{(v_{1},\dots,v_{r}):\prod_{a=1}^{N}R_{a}(v_{1},\dots,v_{r})\not=0\right\}$
is an open set with respect to Zariski topology. In order to show that $\mathbb{U}$ is generic we have to find a $\{v_{1},\dots,v_{r}\}$ such that $\prod_{a}R_{a}(v_{1},\dots,v_{r})\not=0$.
Let $W_{a}=\operatorname{span}\{w_{1,a},\dots,w_{m-r,a}\}$. Notice that each $\operatorname{span}(w_{1,a},\dots,w_{m-r,a}\}$ is a linear subspace of $\mathbb{R}^{m}$ of dimension $m-r$. If $r\geq 1$, each $W_{a}$ is a proper subspace of $\mathbb{R}^{m}$.
A generic $v_{1}\in\mathbb{R}^{m}$ satisfies $v_{1}\not=0$ and $v_{1}\notin\cup_{a=1}^{N}W_{a}$. Replace each $W_{a}$ with $W_{a}^{1}=\operatorname{span}(W_{a},\{v_{1}\})$, subspaces of dimension $\dim(W_{a}^{1})=\dim(W_{a})+1=m-r+1$. If $\dim(W_{a}^{1})<m$, repeat this process inductively and obtain $v_{2},\dots,v_{r}$ until $\dim(W_{a}^{r})=m$. The procedure produces a set of vectors $(v_{1},\dots,v_{r})$ that satisfy the condition $\prod_{a}R_{a}(v_{1},\dots,v_{r})\not=0$. Hence $\mathbb{U}\neq\emptyset$. This ends the proof of 3.1. ∎
Now we apply this lemma to derive the following corollary for our setup:
Corollary 3.2.
Consider the $(N!)^{2p}$ linear maps $L_{\pi_{1},\dots,\pi_{2p}}:{\mathscr{V}}\times{\mathscr{V}}\to\mathbb{R}^{m}$ introduced before. Then for a generic $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d-d_{G}}$,
$ker(\ell)\bigcap\cup_{\pi_{1},...,\pi_{2p}\in S_{N}}Ran(L_{\pi_{1},\dots,\pi_{2p}})=\{0\}$
Proof.
If $m\leq 2d-d_{G}$ then the conclusion is satisfied for any full-rank $\ell$. Therefore assume $m>2d-d_{G}$. A generic linear map $\ell:\mathbb{R}^{m}\to\mathbb{R}^{2d-D_{G}}$ is full-rank. Hence $\dim(\operatorname{Ran}(\ell))=2d-d_{G}$, and thus $\dim(\ker(\ell))=m-2d+d_{G}$. On the other hand, for a generic linear map $\ell$ 3.1 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, and $T=(N!)^{2p}$, implies
$\ker(\ell)\cap\operatorname{Ran}(L_{\pi_{1},\dots\pi_{2p}})=\{0\}$
for every $\pi_{1},\ldots,\pi_{2p}\in S_{N}$. ∎
Lemma 3.3.
Let $\{F_{a}\}_{a=1}^{T}$ be a finite collection of $r$-dimensional subspaces of $\mathbb{R}^{m}$, and $\ell:\mathbb{R}^{m}\to\mathbb{R}^{s}$ be a full-rank linear transformation with $m\geq s$. Let $Q_{a}$ denote the orthogonal projection onto the linear space $F_{a}$ and $Q_{\ell}$ denote the orthogonal projection onto $\ker{\ell}$. Let $c_{a,\ell}=(1-\lVert Q_{a}Q_{\ell}\rVert^{2})^{1/2}$, and $c_{\ell}=\min_{a\in[T]}c_{a,\ell}$. Here $\lVert Q_{a}Q_{\ell}\rVert$ denotes the operator norm of $Q_{a}Q_{\ell}$, i.e., its largest singular value. Set $F=\cup_{a=1}^{T}F_{a}$. Suppose that $\ker(\ell)\cap F=\{0\}$. Then
$\inf_{\begin{subarray}{c}x\in F\\ \lVert x\rVert=1\end{subarray}}\lVert\ell(x)\rVert\geq c_{\ell}\sigma_{s}(\ell),$ (21)
where $\sigma_{s}(\ell)$ is the smallest strictly positive singular value of $\ell$ (it is the $s^{th}$ singular value).
Proof.
Notice that for each $a\in[T]$, the unit sphere of $F_{a}$ is a compact set. Thus
$\inf_{\begin{subarray}{c}x\in F\\ \lVert x\rVert=1\end{subarray}}\lVert\ell(x)\rVert=\min_{\begin{subarray}{c}x\in F\\ \lVert x\rVert=1\end{subarray}}\lVert\ell(x)\rVert=\lVert\ell(y_{\infty})\rVert$
for some $y_{\infty}\in F_{a}\cap S^{1}(\mathbb{R}^{m})$. Let $y_{\infty}=\sum_{k=1}^{m}\gamma_{k}u_{k}$, where $u_{j}$ are the normalized right singular vectors of $\ell$ sorted by singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq\sigma_{q}>\sigma_{s+1}=\cdots=\sigma_{m}=0$. Notice that $\sum_{k=1}^{m}\gamma_{k}^{2}=1$ and $\sum_{k=1}^{s}\gamma_{k}^{2}=1-\lVert Q_{\ell}y_{\infty}\rVert^{2}\geq 1-\lVert Q_{a}Q_{\ell}\rVert^{2}\geq c_{\ell}^{2}$. Thus
$\lVert\ell(y_{\infty})\rVert^{2}$ $=\lVert\sum_{k=1}^{m}\gamma_{k}\ell(u_{k})\rVert^{2}=\lVert\sum_{k=1}^{s}\gamma_{k}\ell(u_{k})\rVert^{2}$
$=\sum_{k=1}^{s}\gamma_{k}^{2}\sigma_{k}^{2}\geq c_{a,\ell}^{2}\sigma_{s}(\ell)^{2}\geq c_{\ell}^{2}\sigma_{s}(\ell)^{2}$
which proves this Lemma. ∎
Proof 1.6.
Assume without loss of generality that $m\geq 2d-d_{G}$.
3.2 shows that, a generic linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d-d_{G}}$ satisfies $\ker(\ell)\cap\operatorname{Ran}(D)=\{0\}$. Thus, if $x,y\in{\mathscr{V}}$ so that $\Psi_{{\textbf{w}},S,\ell}(x)=\Psi_{{\textbf{w}},S,\ell}(y)$ then $\ell(D(x,y))=0$. Therefore $D(x,y)=0$. Since $\hat{\Phi}_{{\textbf{w}},S}$ is injective it follows $x\sim y$. Thus, $\hat{\Psi}_{{\textbf{w}},S,\ell}$ is also injective.
From 1.4 we have that, if the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective then it is also bi-Lipschitz. Let $a\leq b$ denote its bi-Lipschitz constants.
Compositions of two Lipschitz maps is Lipschitz, hence $\Psi_{{\textbf{w}},S,\ell}$ is Lipschitz. Furthermore, an upper Lipschitz constant of $\Psi_{{\textbf{w}},S,\ell}$ is $\lVert\ell\rVert b$, where $\lVert\ell\rVert=\sigma_{1}(\ell)$ is the largest singular value of $\ell$.
Finally from 3.2 and 3.3 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, $T=(N!)^{2p}$, $F_{a}=Ran(L_{\pi_{1},\dots,\pi_{2p}})$ we have that for a generic linear map $\ell$, for all $x,y\in{\mathscr{V}}$,
$\lVert\Psi_{{\textbf{w}},S,\ell}(x)-\Psi_{{\textbf{w}},S,\ell}(y)\rVert$ $=\lVert\ell(D(x,y))\rVert\geq$
$c_{\ell}\sigma_{2d-d_{G}}(\ell)\lVert D(x,y)\rVert$ $\geq c_{\ell}\sigma_{2d-dG}(\ell)a\operatorname{{\mathbf{d}}}([x],[y])$
where $a$ is the lower Lipschitz constant of $\hat{\Phi}_{{\textbf{w}},S}$. Therefore the map $\hat{\Psi}_{{\textbf{w}},S,\ell}$ is bi-Lipschitz with a lower Lipschitz constant $c_{\ell}\sigma_{2d-d_{G}}(\ell)a$. ∎
Remark 3.4.
We proved that if $\hat{\Phi}_{{\textbf{w}},S}$ is injective then for almost any linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d}$, $\hat{\Psi}_{{\textbf{w}},S,\ell}$ is bi-Lipschitz. It remained an open question whether for any such nonlinear embedding $\Psi_{{\textbf{w}},S,\ell}$, injectivity implies bi-Lipschitz. However we settle this question positively into an upcoming joint paper with Matthias Wellershoff. Notice that, in general, if the map $f:X\rightarrow Y$ is bi-Lipschitz and the linear map $\ell:Y\rightarrow\mathbb{R}^{q}$ is so that $\ell\circ f$ is injective, then $\ell\circ f$ may not be bi-Lipschitz. Example: $f:\mathbb{R}\rightarrow\mathbb{R}^{2}$, $f(t)=(t,t^{3})$, $\ell:\mathbb{R}^{2}\rightarrow\mathbb{R}$, $\ell(x,y)=y$.
4 Universality of representation
In previous sections we constructed the embedding $\hat{\Phi}_{{\textbf{w}},S}$ of the quotient space $\hat{{\mathscr{V}}}$ into an Euclidean space $\mathbb{R}^{m}$. In this section we prove 1.7 and 1.8 that show that when $\hat{\Phi}_{{\textbf{w}},S}$ is injective, every continous or Lipschitz map $f:\hat{{\mathscr{V}}}\to H$ factors through $\hat{\Phi}_{{\textbf{w}},S}$.
The proof of 1.7 is based on Kirszbraun extension theorem [kirszbraun1934zusammenziehende] which is re-stated here for the reader’s convenience:
Theorem 4.1 (Kirszbraun extension theorem[kirszbraun1934zusammenziehende]).
Let $E\subset H_{1}$ be an arbitrary subset of a Hilbert space $H_{1}$ and $f:E\mapsto H_{2}$ be a Lipschitz function to another Hilbert space $H_{2}$. Then there exists an extension $F:H_{1}\rightarrow H_{2}$ of $f$ to the entire space $H_{1}$ that has the same Lipschitz constant as the original function $f$.
Proof of 1.7.
  1. 1.
    Let $t:\Phi_{{\textbf{w}},S}({\mathscr{V}})\mapsto H$ be defined by $t(\Phi_{{\textbf{w}},S}(x))=F([x])$. Denote $u=\Phi_{{\textbf{w}},S}(x)$ and $v=\Phi_{{\textbf{w}},S}(y)$. Then,
    $\lVert t(u)-t(v)\rVert=\lVert F([x])-F([y])\rVert\leq\operatorname{Lip}(F)\operatorname{{\mathbf{d}}}([x],[y])\leq\frac{1}{a}\operatorname{Lip}(F)\lVert u-v\rVert.$
    By Kirszbraun extension theorem we have that there exists $T:\mathbb{R}^{m}\to H$, such that
    1. (a)
      $T|_{\Phi_{{\textbf{w}},S}({\mathscr{V}})}=t$
    2. (b)
      $\operatorname{Lip}(T)=\operatorname{Lip}(t)$
    Therefore, $F=T\circ\Phi_{{\textbf{w}},S}$ and $\operatorname{Lip}(T)\leq\frac{1}{a}\operatorname{Lip}(F)$.
  2. 2.
    Part 2 is straightforward. Let $x,y\in{\mathscr{V}}$ then
    $\lVert F(x)-F(y)\rVert$ $=\lVert T\circ\Phi_{{\textbf{w}},S}(x)-T\circ\Phi_{{\textbf{w}},S}(y)\rVert$
    $\leq\operatorname{Lip}(T)\lVert\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)\rVert$
    $\leq\operatorname{Lip}(T)a\lVert x-y\rVert.$
The second universality result 1.8 applies to the class of continuous functions (instead of Lipschitz functions).
The proof of 1.8 follows from the following extension of Tietze’s theorem [urysohn1925machtigkeit, dugundji1951extension].
Theorem 4.2 (Dugundji-Tietze [dugundji1951extension]).
Let $X$ be a metric space and $A$ a closed subset of $X$. Let $L$ be a locally convex topological vector space. Given $f:A\mapsto L$ a continuous map, there exists a continuous extension $F:X\mapsto L$ such as $F(X)$ is a subset of the convex hull of $f(A)$.
Proof of 1.8.
Let $S=\Phi_{{\textbf{w}},S}({\mathscr{V}})\subset\mathbb{R}^{m}$. Note that $S$ is a closed set as $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz and $\hat{{\mathscr{V}}}$ is complete. Let $t:S\mapsto L$ be defined by $t(\Phi_{{\textbf{w}},S}(x))=F(x)$, for all $x\in{\mathscr{V}}$. Note $t$ is continuous since $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz on $\hat{{\mathscr{V}}}$. By 4.2 there exists a continuous extension $T:\mathbb{R}^{m}\rightarrow L$ of $t$ that satisfies the convex hull property.
[…]

### v4

@@
-bibliography.bib
+G-Invariant Representations using Coorbits: Bi-Lipschitz Properties
 Abstract
@@
 This embedding $\Psi$ should also separate data orbits and satisfy certain stability conditions to ensure that small perturbations of the input do not significantly impact the predictions.
-We worked in the orbit separation problem in [balan2023inj].
+We worked in the orbit separation problem in [11].
 In this paper we focus on the stability problem.
-This problem is an instance of invariant machine learning. [mixon2022injectivity, aslan2023group, dym2022low, bronstein, maron2018invariant, DUFRESNE20091979, cahill2023bilipschitz].
-The most common group actions in invariant machine learning are permutations [sannai2020universal, cahill22, balan2022permutation] reflections [mixon2022max] and translations [Cahill19].
-A related, by somewhat different problem is the case of equivariant embeddings [lipman2022, maron2018invariant, villar2021scalars, villar2022dimensionless, blum2022equivariant].
-Our work extends and unifies two prior approaches: the max filter embedding introduced in [cahill22, mixon2022injectivity], and the permutation invariant representation introduced in [balan2022permutation].
-The construction of permutation invariant embeddings is closely connected to the phase retrieval problem [balan2004signal, bcmn] that has a large body of results.
-For instance, [balan16] provides exact estimates for both the upper and lower Lipschitz bounds, addressing both the real and complex cases of the phase retrieval problem.
-In [Balan2023Rel] we establish an isometric identification of the real phase retrieval problem to $S_{2}$-invariant representations.
-A completely different approach is considered in [eriksson2018quantitative].
+This problem is an instance of invariant machine learning. [30, 5, 22, 16, 28, 20, 17, 19].
+The most common group actions in invariant machine learning are permutations [33, 18, 10] reflections [29] and translations [25].
+A related, by somewhat different problem is the case of equivariant embeddings [31, 28, 37, 36, 15].
+The phase retrieval that was introduced in [9] is an instance of this setup, when the compact group is the torus. [8, 6, 7, 3, 24, 2].
+Our work extends and unifies two prior approaches: the max filter embedding introduced in [18, 30], and the permutation invariant representation introduced in [10].
+[4] analyzes the stability of generalized phase retrieval problems under the action of compact groups, providing conditions under which stable recovery is possible from invariant measurements.
+The construction of permutation invariant embeddings is closely connected to the phase retrieval problem [9, 1] that has a large body of results.
+For instance, [6] provides exact estimates for both the upper and lower Lipschitz bounds, addressing both the real and complex cases of the phase retrieval problem.
+In [13] we establish an isometric identification of the real phase retrieval problem to $S_{2}$-invariant representations.
+A completely different approach is considered in [23].
 There it is proved that for any discrete group $G$ of isometries acting on $\mathbb{R}^{d}$, it is possible to construct a bi-Lipschitz map from $\mathbb{R}^{d}/G$ to $\mathbb{R}^{N}$, where the distortion depends solely on $d$.
-Since the first draft of this paper was placed on arxiv, the authors of [qaddura2024stable] extended the construction of this paper and proved that for certain compact groups, given enough generic templates, the coorbit filter bank (as defined there) is injective and bi-Lipschitz.
+Since the first draft of this paper was placed on arxiv, the authors of [32] extended the construction of this paper and proved that for certain compact groups, given enough generic templates, the coorbit filter bank (as defined there) is injective and bi-Lipschitz.
 In this paper we construct an Euclidean embedding that is globally bi-Lipschitz and can be implemented relatively easy.
@@
 1.1 Notation
-In this paper we use the same notation as in [balan2023inj].
+In this paper we use the same notation as in [11].
 Let $({\mathscr{V}},{\langle\cdot,\cdot\rangle})$ be a $d$-dimensional real vector space with scalar product, and $d\geq 2$.
@@
 We denote by $[x]$ the orbit of vector $x$, i.e.
-$[x]=\{U_{g}x:g\in G\}$.
+$x]=\{U_{g}x:g\in G\}$.
 The natural metric, $\operatorname{{\mathbf{d}}}:\hat{{\mathscr{V}}}\times\hat{{\mathscr{V}}}\to\mathbb{R}$, is defined by
@@
 for some permutation $\sigma:\{1,\ldots,r\}\rightarrow\{1,\ldots,r\}$.
-For an integer $p\in\mathbb{N}$, we denote $[p]=\{1,2,\ldots,p\}$.
+For an integer $p\in\mathbb{N}$, we denote $p]=\{1,2,\ldots,p\}$.
 For a set $S$, $|S|$ denotes its cardinal.
@@
 $\Phi_{{\textbf{w}},S}:{\mathscr{V}}\rightarrow\mathbb{R}^{m}\leavevmode\nobreak\ ,\leavevmode\nobreak\ \Phi_{{\textbf{w}},S}(x)=[\{\Phi_{w_{1},j}(x)\}_{j\in S_{1}},\dots,\{\Phi_{w_{p},j}(x)\}_{j\in S_{p}}]\in\mathbb{R}^{m}.$ (2)
-$[y]$$[x]$$\hat{{\mathscr{V}}}$$\mathbb{R}^{m}$$\mathbb{R}^{q}$$\Phi_{{\textbf{w}},S}$$\Phi_{{\textbf{w}},S}$$\ell$$\ell$$\Phi_{{\textbf{w}},S}([x])$$\Phi_{{\textbf{w}},S}([y])$$\ell\circ\Phi_{{\textbf{w}},S}([y])$$\ell\circ\Phi_{{\textbf{w}},S}([x])$
+[…]
 Figure 1: Proposed embedding scheme
@@
 In this paper we focus on stability and universality properties of maps $\Phi_{{\textbf{w}},S}$ and $\Psi_{{\textbf{w}},S,\ell}$.
-Informally, our main results, 1.4, 1.6, 1.7 and 1.8, state that: (1) "injectivity" implies "(bi-Lipschitz) stability", (2) stable bi-Lipschitz embedding can be achieved into an Euclidean space of dimension at most twice the dimension of the input space; and (3) any continuous or even Lipschitz map factors through $\Phi_{{\textbf{w}},S}$.
+Informally, our main results, ˜1.4, ˜1.6, ˜1.7 and ˜1.8, state that: (1) "injectivity" implies "(bi-Lipschitz) stability", (2) stable bi-Lipschitz embedding can be achieved into an Euclidean space of dimension at most twice the dimension of the input space; and (3) any continuous or even Lipschitz map factors through $\Phi_{{\textbf{w}},S}$.
 For the rest of the paper we shall use interchangeably $\Phi_{i,j}$ instead of $\Phi_{w_{i},j}$.
@@
 Theorem 1.8.
-Assume that, for fixed $S\in[N]\times[p]$ and ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$, the map $\hat{\Phi}_{{\textbf{w}},S}:{\mathscr{V}}\to\mathbb{R}^{m}$ is bi-Lipschitz on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$.
-Let $L$ denote a locally convex topological vector space.
-For every continuous map $F:\hat{{\mathscr{V}}}\to L$ there exists a continuous $T:\mathbb{R}^{m}\to L$ such that $F=T\circ\Phi_{{\textbf{w}},S}$.
-Furthermore $T(\mathbb{R}^{m})$ is included in the convex hull of $F(\hat{{\mathscr{V}}})$.
+[…]
+Remark 1.9.
+In [14] we generalize ˜1.4 for the case where our embedding has the form $a\circ\Phi_{{\textbf{w}},S}(x)$, where $a$ is a linear map.
+Remark 1.10.
+In [12] we study under which conditions the map $\hat{\Phi}_{{\textbf{w}},S}:\hat{{\mathscr{V}}}\to\mathbb{R}^{m}$ is injective on the quotient space $\hat{{\mathscr{V}}}={\mathscr{V}}/G$.
 2 Stability of Embedding
 Suppose that for ${\textbf{w}}=(w_{1},\dots,w_{p})\in{\mathscr{V}}^{p}$ and $S\subset[N]\times[p]$ the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective.
-In this case, we claim the map $\hat{\Phi}_{{\textbf{w}},S}$ is also bi-Lipschitz, (1.4).
-An estimate of the upper Lipschitz constant $b$ is given in 2.1.
-1.5 follows from 1.4.
-The proof of 1.4 is contained in the next two subsections.
+In this case, we claim the map $\hat{\Phi}_{{\textbf{w}},S}$ is also bi-Lipschitz, (˜1.4).
+An estimate of the upper Lipschitz constant $b$ is given in ˜2.1.
+˜1.5 follows from ˜1.4.
+The proof of ˜1.4 is contained in the next two subsections.
 2.1 Upper Lipschitz bound
@@
 By the fundamental theorem of calculus we get
-[…]
+$\Phi_{i,j}(x)-\Phi_{i,j}(y)=\int_{0}^{1}\frac{d}{dt}\Phi_{i,j}((1-t)x+ty)dt.$
 Therefore,
-[…]
+$\Phi_{i,j}(y)-\Phi_{i,j}(x)=\int_{0}^{1}{\langle y-x,U_{g_{j_{t}}}w_{i}\rangle}dt$
 so
-[…]
+$\lVert\operatorname{\downarrow}\{\Phi_{i,j}(x)\}_{j\in S_{i}}-\operatorname{\downarrow}\{\Phi_{i,j}(y)\}_{j\in S_{i}}\rVert$ $\leq\int_{0}^{1}(\sum_{j\in S_{i}}{\langle y-x,U_{g_{j_{t}}}w_{i}\rangle}^{2})^{1/2}dt$
 $\leq\sqrt{B_{i}}\lVert x-y\rVert$
@@
 Consider also the map
-$\Delta^{i,j}(x)=\begin{cases}\min_{g\notin L^{i,j}(x)}(|{\langle U_{g}w_{i},x\rangle}-\Phi_{i,j}(x)|)\frac{1}{\lVert w_{i}\rVert},\leavevmode\nobreak\ \text{ if }L^{i,j}(x)\not=G\\ \frac{\lVert x\rVert}{\lVert w_{i}\rVert},\leavevmode\nobreak\ \text{ if }L^{i,j}(x)=G.\\ \end{cases}$ (5)
+[…]
 Lemma 2.2.
@@
 $\left|\{g\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{g}w_{i},x\rangle}<\Phi_{i,j}(x)\}\right|>N-j$
-then there exist at least $N-j+1$, distinct elements of group $G$, say $\{h_{1},\dots h_{j}\}$ such that ${\langle U_{h_{k}}w_{i},x\rangle}<{\langle U_{g}w_{i},x\rangle}\leavevmode\nobreak\ \forall k\in[N-j+1]$, which is also a contradiction.
+then there exist at least Nj+1N-j+1, distinct elements of group $G$, say $\{h_{1},\dots h_{j}\}$ such that ${\langle U_{h_{k}}w_{i},x\rangle}<{\langle U_{g}w_{i},x\rangle}\leavevmode\nobreak\ \forall k\in[N-j+1]$, which is also a contradiction.
 (b) If $\Phi_{i,j}(x)=\Phi_{i,j+1}(x)$, a group element $g$ achieves $\Phi_{i,j}(x)$ if and only if, it also achieves $\Phi_{i,j+1}(x)$, therefore $L^{i,j}(x)=L^{i,j+1}(x)$.
@@
 $\cup_{k\leq j}L^{i,k}(x)\subset\{h\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{h}w_{i},x+y\rangle}>\Phi_{i,j}(x+y)\}.$
-But the set $\cup_{k\leq j}L^{i,k}(x)$ contains at least $j$ elements (since each $L^{i,j}(x)$ is non-empty) and so we derived a contradiction with 2.2(a) Equation 6.
+But the set $\cup_{k\leq j}L^{i,k}(x)$ contains at least $j$ elements (since each $L^{i,j}(x)$ is non-empty) and so we derived a contradiction with ˜2.2(a) Equation˜6.
 Lemma 2.4.
@@
 Then $\Phi_{i,q}(x+c_{1}y)={\langle U_{g_{1}}w_{i},x+c_{1}y\rangle}<\Phi_{i,j}(x+c_{1}y)$.
-By 2.2 (b)(ii),
+By ˜2.2 (b)(ii),
 $\left|\cup_{r\leq j}L^{i,r}(x+c_{1}y)\right|=q-1\geq j,$
-and $g_{1}\not\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$. On the other hand, from 2.2 (a), Equation 6,
+and $g_{1}\not\in\cup_{r\leq j}L^{i,r}(x+c_{1}y)$. On the other hand, from ˜2.2 (a), Equation˜6,
 $\left|\{h\in G\leavevmode\nobreak\ ,\leavevmode\nobreak\ {\langle U_{h}w_{i},x+c_{2}y\rangle}>\Phi_{i,j}(x+c_{2}y)\}\right|\leq j-1$
@@
 ${\langle U_{h}w_{i},x+c_{2}y\rangle}\leq\Phi_{i,j}(x+c_{2}y)={\langle U_{g_{1}}w_{i},x+c_{2}y\rangle}.$ (8)
-On the other hand, by 2.3, $g_{1}\in L^{i,j}(x)$.
+On the other hand, by ˜2.3, $g_{1}\in L^{i,j}(x)$.
 But if ${\langle U_{h}w_{i},x\rangle}-{\langle U_{g_{1}}w_{i},x\rangle}>0$ then
@@
 1.
-Note that by norm conditions on $z_{r}$, 2.3 and 2.4, if $g_{1},g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}\leavevmode\nobreak\ \forall l\in[k]$.
+Note that by norm conditions on $z_{r}$, ˜2.3 and ˜2.4, if $g_{1},g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$ then ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{l}z_{r}\rangle}\leavevmode\nobreak\ \forall l\in[k]$.
 Starting with $l=1$ and proceeding recursively we get ${\langle U_{g_{1}}w_{i},z_{r}\rangle}={\langle U_{g_{2}}w_{i},z_{r}\rangle}\leavevmode\nobreak\ \forall r\in[k]$.
@@
 2.
-$\subset$". From 2.3 we have that $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$.
-$\supset$". Let $g_{1}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$. Take any $g_{2}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. By 2.3, $g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$, but from part 1 of this Lemma,
+$\subset"$. From ˜2.3 we have that $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$.
+$\supset"$. Let $g_{1}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$. Take any $g_{2}\in L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$. By ˜2.3, $g_{2}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$, but from part 1 of this Lemma,
 ${\langle U_{g_{1}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}={\langle U_{g_{2}}w_{i},\sum_{r=1}^{k}a_{r}z_{r}\rangle}.$
@@
 3.
-Case 1. Suppose $L^{i,j}(\sum_{r=1}^{k}z_{r})=G$. From 2.3 we know that
+Case 1. Suppose $L^{i,j}(\sum_{r=1}^{k}z_{r})=G$. From ˜2.3 we know that
 $G=L^{i,j}(\sum_{r=1}^{k}z_{r})\subset L^{i,j}(\sum_{r=1}^{k-1}z_{r})\subset\dots\subset L^{i,j}(z_{1}).$
@@
 $>0.$
-But from 2.2 $|\cup_{l\geq j}L^{i,l}L^{i,l}(\sum_{r=1}^{k}z_{r})|\geq N-j+1$. So $g_{1}\in\cup_{l=1}^{j}L^{i,l}(\sum_{r=1}^{k}z_{r})$. Therefore,
+But from ˜2.2 $|\cup_{l\geq j}L^{i,l}L^{i,l}(\sum_{r=1}^{k}z_{r})|\geq N-j+1$. So $g_{1}\in\cup_{l=1}^{j}L^{i,l}(\sum_{r=1}^{k}z_{r})$. Therefore,
 $\cup_{l\leq j}L^{i,l}(\sum_{r=1}^{k}z_{r})\subset\cup_{l\leq j}L^{i,l}(\sum_{r=1}^{k}a_{r}z_{r}).$
@@
 $\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}=\left(L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\right)^{c}.$
-First notice that by 2.3 we have that $L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$ and $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$.
-From 2.6 part (2) we have that $L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$.
+First notice that by ˜2.3 we have that $L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}z_{r})$ and $L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\subset L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$.
+From ˜2.6 part (2) we have that $L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$.
 Take $g\in L^{i,j}(\sum_{r=1}^{k}z_{r}+e)$ and $h\in\left(L^{i,j}(\sum_{r=1}^{k}z_{r}+e)\right)^{c}$.
@@
 $-\left|{\langle U_{g}w_{i},\sum_{r=1}^{k}(a_{r}-1)z_{r}+e\rangle}\right|\geq$
-[…]
+$\geq\lVert w_{i}\rVert\left[\Delta^{i,j}(\sum_{r=1}^{k}z_{r})-2\left(\sum_{r=1}^{k}|a_{r}-1|\lVert z_{r}\rVert+\lVert e\rVert\right)\right]>0\hskip 56.9055pt(*)$
 Hence again $h\in\left(L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r}+e)\right)^{c}$.
@@
   • Next we construct inductively a sequence of non-zero vectors $z_{2},z_{3},...,z_{k}$ so that the local lower Lipschitz constant vanishes in a convex set of the form $\{\sum_{r=1}^{k}a_{r}z_{r}\leavevmode\nobreak\ ,\leavevmode\nobreak\ |a_{r}-1|<\delta\}$ for some $\delta>0$ small enough, and where sets $L^{i,j}$ remain constant.
-These steps are depicted in Figure 3 that suggests a "robotic arm" procedure (this name was suggested by the authors of [qaddura2024stable]).
+These steps are depicted in Figure 3 that suggests a "robotic arm" procedure (this name was suggested by the authors of [32]).
   • For $k=d$ this construction defines a non-empty open set where the local lower Lipschitz constant vanishes and $L^{i,j}$ remain constants.
@@
 ∎
-$z_{1}$$x$$y$$z$$x_{n}$$y_{n}$
+[…]
 Figure 2: Local analysis of sequences $x_{n}$ and $y_{n}$ converging to $z_{1}$.
@@
 Remark 2.10.
-Applying Lemma 2.9 to two sequences $(x_{k})_{k}$ and $(y_{k})_{k}$ that satisfy (14-17) in 2.8, it follows that $d(x_{k},z_{1})=\lVert x_{k}-z_{1}\rVert$ and $d(y_{k},z_{1})=\lVert y_{k}-z_{1}\rVert$ for $k$ large enough.
+Applying Lemma 2.9 to two sequences $(x_{k})_{k}$ and $(y_{k})_{k}$ that satisfy (14-17) in ˜2.8, it follows that $d(x_{k},z_{1})=\lVert x_{k}-z_{1}\rVert$ and $d(y_{k},z_{1})=\lVert y_{k}-z_{1}\rVert$ for $k$ large enough.
 Hence alignment must occur from some rank on.
 Lemma 2.11.
-For fixed $i\in[p]$, $j\in S_{i}$ and two sequences $(x_{n})_{n}$, $(y_{n})_{n}$ produced by 2.8, we denote by $g_{1,n,i,j}$ the group elements that achieves $\Phi_{i,j}(x_{n})$ and by $g_{2,n,i,j}$ the group element that achieves the $\Phi_{i,j}^{j}(y_{n})$.
+For fixed $i\in[p]$, $j\in S_{i}$ and two sequences $(x_{n})_{n}$, $(y_{n})_{n}$ produced by ˜2.8, we denote by $g_{1,n,i,j}$ the group elements that achieves $\Phi_{i,j}(x_{n})$ and by $g_{2,n,i,j}$ the group element that achieves the $\Phi_{i,j}^{j}(y_{n})$.
 That is $\Phi_{i,j}(x_{n})={\langle U_{g_{1,n,i,j}}w_{i},x_{n}\rangle}$ and $\Phi_{i,j}(y_{n})={\langle U_{g_{2,n,i,j}}w_{i},y_{n}\rangle}$.
@@
 So by induction after $\sum_{i\in[p]}m_{i}=m$ steps we construct a subsequence of $(x_{n})_{n}$ lets call it $(x_{n_{m}})_{m}$ such that $g_{1,i,j,n_{m}}=g_{1,i,j}$ for every $i\in[p],j\in S_{i}$.
-Starting from sequence $(y_{1,n_{m}})_{m}$ we repeat the same procedure concluding in a subsequence $(y_{1,n_{r}})_{r}$ such that $g_{2,i,j,n_{r}}=g_{2,i,j}$ for every $r\in\mathbb{N},i\in[p],j\in S_{i}$ Notice that sequences $(x_{n_{r}})_{r}$ and $(y_{n_{r}})_{r}$ that from now on we will call them $(x_{n})_{n}$ and $(y_{n})_{n}$ for easiness of notation, satisfy the assumptions of lemma.
+Starting from sequence $(y_{1,n_{m}})_{m}$ we repeat the same procedure concluding in a subsequence $(y_{1,n_{r}})_{r}$ such that $g_{2,i,j,n_{r}}=g_{2,i,j}$ for every $r\in\mathbb{N},i\in[p],j\in S_{i}$ .
+Notice that sequences $(x_{n_{r}})_{r}$ and $(y_{n_{r}})_{r}$ that from now on we will call them $(x_{n})_{n}$ and $(y_{n})_{n}$ for easiness of notation, satisfy the assumptions of lemma.
 ∎
@@
 $+$ ${\langle w_{i},U_{g_{1,i,j}^{-1}}u_{n}-U_{g_{2,i,j}^{-1}}v_{n}\rangle}|^{2}.$
-This sequence converge to 00, as $k\to\infty$ while also $u_{n},v_{n}\to 0$.
+This sequence converge to 0, as $k\to\infty$ while also $u_{n},v_{n}\to 0$.
 So we conclude that for each $i\in[p]$ and $j\in S_{i}$, ${\langle U_{g_{1,i,j}}w_{i}-U_{g_{2,i,j}}w_{i},z_{1}\rangle}=0$.
@@
 where $\lVert u_{n}\rVert,\lVert v_{n}\rVert\to 0$, so for large enough $n$ we have that $\lVert u_{n}\rVert,\lVert v_{n}\rVert\leq\frac{1}{4}\rho_{0}(z_{1})$.
-Recall that from 2.9, we conclude that exists $N_{0}\in\mathbb{N}$, such that $\lVert u_{n}-v_{n}\rVert\leq\lVert U_{g}u_{n}-v_{n}\rVert$ for all $g\in H(z_{1})$ and $k\geq N_{0}$.
+Recall that from ˜2.9, we conclude that exists $N_{0}\in\mathbb{N}$, such that $\lVert u_{n}-v_{n}\rVert\leq\lVert U_{g}u_{n}-v_{n}\rVert$ for all $g\in H(z_{1})$ and $k\geq N_{0}$.
 Lemma 2.12.
@@
       $\lVert v_{n}\rVert=\lVert\tilde{y}_{n}-\sum_{r=1}^{k}a_{r}z_{r}\rVert<\frac{1}{16}\Delta(\sum_{r=1}^{k}a_{r}z_{r})$.
-    Thus from part (3) of 2.6, 2.3 and part (2) of 2.6 we have that for any $n\geq M_{0}$ and $(i,j)\in S$
+    Thus from part (3) of ˜2.6, ˜2.3 and part (2) of ˜2.6 we have that for any $n\geq M_{0}$ and $(i,j)\in S$
     $L^{i,j}(\tilde{x}_{n})=L^{i,j}(x_{n})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}a_{r}z_{r})$
@@
     Let two sequences $(x_{n})_{n},(y_{n})_{n}$ that both converge to $\sum_{r=1}^{k}z_{r}$, and achieve lower Lipschitz bound zero for map $\Phi_{{\textbf{w}},S}$.
-We align sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ to satisfy the properties of 2.8.
+We align sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ to satisfy the properties of ˜2.8.
 We denote by $a_{n}=P_{E_{k}}x_{n}$ and $b_{n}=P_{E_{k}}y_{n}$ the orthogonal projections of the sequences $(x_{n})_{n}$ and $(y_{n})_{n}$ respectively, on the linear subspace $E_{k}=\operatorname{span}\{z_{1},\dots,z_{k}\}^{\bot}$.
@@
 Assuming otherwise, there are two sequences of vectors $x_{n}=\sum_{r=1}^{k}c_{r,n}z_{r}$ and $y_{n}=\sum_{r=1}^{k}d_{r,n}z_{r}$, where $\lim_{n\to\infty}c_{r,n}=\lim_{n\to\infty}d_{r,n}=1,\leavevmode\nobreak\ \forall r\in[k]$ that achieve lower Lipschitz bound zero.
-Recall that from part (2) of 2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
+Recall that from part (2) of ˜2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
     $L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{k}d_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{k}z_{r}).$
     Then, for $g_{i,j}\in L^{i,j}(\sum_{r=1}^{k}z_{r})$,
-    00 0 $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
+    $0$ $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(x_{n})-\Phi_{{\textbf{w}},S}(y_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
     $=\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},x_{n}-y_{n}\rangle}|^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
@@
 In this case
-    $\Phi_{{\textbf{w}},S}(\sum_{r=1}^{k}z_{r})=\Phi_{{\textbf{w}},S}(\sum_{r=1}^{k}z_{r}+\epsilon\tilde{z})$
+[…]
     which contradict the injectivity property. This establishes Claim 1.
@@
       $\lVert e_{n}\rVert<\frac{1}{16k}\Delta(\sum_{r=1}^{k}z_{r})$
-    From 2.6 part (1),
+    From ˜2.6 part (1),
     $\Delta(\sum_{r=1}^{k}c_{r,n}z_{r})\geq\frac{1}{4}\Delta(\sum_{r=1}^{k}z_{r}).$
@@
     So, for any $(i,j)\in S$
-    $L^{i,j}(x_{n})$ $=L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r}+a_{n})=L^{i,j}(\sum_{r=1}^{k}z_{r}+a_{n})$
+    $\hskip-28.45274ptL^{i,j}(x_{n})$ $=L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r}+a_{n})=L^{i,j}(\sum_{r=1}^{k}z_{r}+a_{n})$
     $=L^{i,j}(\sum_{r=1}^{k}z_{r}+s_{n}a_{n})=L^{i,j}(\tilde{x}_{n})\subset L^{i,j}(\sum_{r=1}^{k}z_{r})=L^{i,j}(\sum_{r=1}^{k}c_{r,n}z_{r}).$
-    Where the second equality comes from 2.6 part 3, third equality from 2.4 the fifth inclusion from 2.3, and the last equality from 2.6 part 2.
+    Where the second equality comes from ˜2.6 part 3, third equality from ˜2.4 the fifth inclusion from ˜2.3, and the last equality from ˜2.6 part 2.
     Similarly,
-    $L^{i,j}(y_{n})$ $=L^{i,j}(\sum_{r=1}^{k}d_{r,n}z_{r}+b_{n})=L^{i,j}(\sum_{r=1}^{k}(1+d_{r,n}-c_{r,n})z_{r}+b_{n})$
+    $\hskip-28.45274ptL^{i,j}(y_{n})$ $=L^{i,j}(\sum_{r=1}^{k}d_{r,n}z_{r}+b_{n})=L^{i,j}(\sum_{r=1}^{k}(1+d_{r,n}-c_{r,n})z_{r}+b_{n})$
     $=L^{i,j}(\sum_{r=1}^{k}z_{r}+e_{n})=L^{i,j}(\sum_{r=1}^{k}z_{r}+s_{n}e_{n})=L^{i,j}(\tilde{y}_{n})$
@@
     Claim 3: The two vectors $\alpha$ and $\delta$ are equal, $\alpha=\delta$.
-    We prove this claim by contradiction. Assume that $\alpha\neq\delta$. From 2.9, $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$
+    We prove this claim by contradiction. Assume that $\alpha\neq\delta$. From ˜2.9, $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$
     $\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert.$
     Therefore,
-    $00<\lVert a-\delta\rVert=\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert=$ 0 .
+    $0<\lVert a-\delta\rVert=\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}e_{n}\rVert\leq\lim_{n\to\infty}\lVert s_{n}a_{n}-s_{n}U_{g_{1}}e_{n}\rVert=0.$
     We conclude that $\alpha=\delta\neq 0$.
@@
 $\frac{1}{16}\min\left(\lVert z_{k}\rVert,\Delta(\sum_{r=1}^{k}z_{r}),\rho_{0}(\sum_{r=1}^{k}z_{r})\right)$.
-Now we can complete the proof of 1.4.
+Now we can complete the proof of ˜1.4.
 $z_{1}$$z_{2}$
@@
 Proof.
-Starting from vector $z_{1}$ and the sequences $(x_{n})_{n}$, $(y_{n})_{n}$ observed in 2.1 after $d-1$ steps of algorithmic construction of part (2) of 2.12 we get $d$ non-zero vectors $\{z_{1},\dots,z_{d}\}$ and a pair of sequences $(\tilde{x}_{n})_{n}$, $(\tilde{y}_{n})_{n}$ such that
+Starting from vector $z_{1}$ and the sequences $(x_{n})_{n}$, $(y_{n})_{n}$ observed in ˜2.1 after d1d-1 steps of algorithmic construction of part (2) of ˜2.12 we get $d$ non-zero vectors $\{z_{1},\dots,z_{d}\}$ and a pair of sequences $(\tilde{x}_{n})_{n}$, $(\tilde{y}_{n})_{n}$ such that
 (i) ${\langle z_{i},z_{j}\rangle}=0,\leavevmode\nobreak\ \forall i,j\in[d],i\neq j$;
@@
 Let $\tilde{x}_{n}=\sum_{r=1}^{d}l_{r,n}z_{r}$ and $\tilde{y}_{n}=\sum_{r=1}^{d}t_{r,n}z_{r}$. Notice that $\lim_{n\to\infty}l_{r,n}=\lim_{n\to\infty}t_{r,n}=1,\leavevmode\nobreak\ \forall r\in[d]$.
-Recall that from part (2) of 2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
+Recall that from part (2) of ˜2.6 we have that $\exists M_{0}\in\mathbb{N}$ such that $\forall n\geq M_{0}$ and $(i,j)\in S$
 $L^{i,j}(\sum_{r=1}^{d}l_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{d}t_{r,n}z_{r})=L^{i,j}(\sum_{r=1}^{d}z_{r}).$
 Then, for $g_{i,j}\in L^{i,j}(\sum_{r=1}^{d}z_{r})$,
-00 0 $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
+$0$ $=\lim_{n\to\infty}\frac{\lVert\Phi_{{\textbf{w}},S}(\tilde{x}_{n})-\Phi_{{\textbf{w}},S}(\tilde{y}_{n})\rVert^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
 $=\lim_{n\to\infty}\frac{\sum_{i=1}^{p}\sum_{j\in S_{i}}|{\langle U_{g_{i,j}}w_{i},\tilde{x}_{n}-\tilde{y}_{n}\rangle}|^{2}}{\operatorname{{\mathbf{d}}}([x_{n}],[y_{n}])^{2}}$
@@
 In this case
-$\Phi_{{\textbf{w}},S}(\sum_{r=1}^{d}z_{r})=\Phi_{{\textbf{w}},S}(\sum_{r=1}^{d}z_{r}+\epsilon\tilde{z})$
-which contradict the injectivity property. 1.4 is now proved.
+[…]
+which contradict the injectivity property. ˜1.4 is now proved.
 3 Dimension reduction using linear maps
@@
 The idea of using dimension reduction linear maps goes back many years.
-The famous Johnson-Lindenstrauss Lemma [johnson1986extensions] provides a nearly isometric projection for finite metric spaces.
-More recently, [Cahill19] shows that a dimension reduction linear map preserves Lipschitz properties for translation invariant polynomial embeddings.
-The first result that combines sorted co-orbits with linear maps was shown in [balan2022permutation] in the case of the group of permutations $G=S_{n}$ acting by left multiplication on $n\times d$ real matrices, ${\mathscr{V}}=\mathbb{R}^{n\times d}$.
-In that particular case, the dimension of the intermediary co-orbit space was $n(1+(d-1)n!)$ while the final target space was shown to be $2nd=2\,dim({\mathscr{V}})$.
-The authors of [dym2022low] show that intermediate embedding can be realized in dimension $(2nd+1)n$ instead of $n(1+(d-1)n!)$ with the final target space of dimension $12nd+$ 1.
-A more careful analysis of such an embedding by Matthias Wellershoff [Matthias2023a] proved that $12nd+$ 1 can be replaced by $2nd-d$.
+The famous Johnson-Lindenstrauss Lemma [26] provides a nearly isometric projection for finite metric spaces.
+More recently, [25] shows that a dimension reduction linear map preserves Lipschitz properties for translation invariant polynomial embeddings.
+The first result that combines sorted co-orbits with linear maps was shown in [10] in the case of the group of permutations $G=S_{n}$ acting by left multiplication on $n\times d$ real matrices, ${\mathscr{V}}=\mathbb{R}^{n\times d}$.
+In that particular case, the dimension of the intermediary co-orbit space was n(1+(d1)n!)n(1+(d-1)n!) while the final target space was shown to be $2nd=2\,dim({\mathscr{V}})$.
+The authors of [22] show that intermediate embedding can be realized in dimension $(2nd+1)n$ instead of n(1+(d1)n!)n(1+(d-1)n!) with the final target space of dimension $2nd+1$.
+A more careful analysis of such an embedding by Matthias Wellershoff [38] proved that $2nd+1$ can be replaced by 2ndd2nd-d.
 In the Ph.D.
-thesis [tsoukanis2024g], one of the authors showed that the decrease of dimension from $2nd$ to $2nd-d$ represents the dimension of a certain joint eigenspace of this representation.
-The exact statement of our dimension reduction result is included in 1.6.
-To prove 1.6 we need first to construct a few objects.
+thesis [34], one of the authors showed that the decrease of dimension from $2nd$ to 2ndd2nd-d represents the dimension of a certain joint eigenspace of this representation.
+The exact statement of our dimension reduction result is included in ˜1.6.
+To prove ˜1.6 we need first to construct a few objects.
 Let $D:{\mathscr{V}}\times{\mathscr{V}}\to\mathbb{R}^{m}$ be the nonlinear map $D(x,y)=\Phi_{{\textbf{w}},S}(x)-\Phi_{{\textbf{w}},S}(y)$.
@@
 Let $W_{a}=\operatorname{span}\{w_{1,a},\dots,w_{m-r,a}\}$.
-Notice that each $\operatorname{span}(w_{1,a},\dots,w_{m-r,a}\}$ is a linear subspace of $\mathbb{R}^{m}$ of dimension $m-r$.
+Notice that each $\operatorname{span}(w_{1,a},\dots,w_{m-r,a}\}$ is a linear subspace of $\mathbb{R}^{m}$ of dimension mrm-r.
 If $r\geq 1$, each $W_{a}$ is a proper subspace of $\mathbb{R}^{m}$.
@@
 Hence $\mathbb{U}\neq\emptyset$.
-This ends the proof of 3.1.
+This ends the proof of ˜3.1.
 ∎
@@
 Hence $\dim(\operatorname{Ran}(\ell))=2d-d_{G}$, and thus $\dim(\ker(\ell))=m-2d+d_{G}$.
-On the other hand, for a generic linear map $\ell$ 3.1 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, and $T=(N!)^{2p}$, implies
+On the other hand, for a generic linear map $\ell$ ˜3.1 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, and $T=(N!)^{2p}$, implies
 $\ker(\ell)\cap\operatorname{Ran}(L_{\pi_{1},\dots\pi_{2p}})=\{0\}$
@@
 which proves this Lemma. ∎
-Proof 1.6.
+Proof ˜1.6.
 Assume without loss of generality that $m\geq 2d-d_{G}$.
-3.2 shows that, a generic linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d-d_{G}}$ satisfies $\ker(\ell)\cap\operatorname{Ran}(D)=\{0\}$.
+˜3.2 shows that, a generic linear map $\ell:\mathbb{R}^{m}\rightarrow\mathbb{R}^{2d-d_{G}}$ satisfies $\ker(\ell)\cap\operatorname{Ran}(D)=\{0\}$.
 Thus, if $x,y\in{\mathscr{V}}$ so that $\Psi_{{\textbf{w}},S,\ell}(x)=\Psi_{{\textbf{w}},S,\ell}(y)$ then $\ell(D(x,y))=0$.
@@
 Thus, $\hat{\Psi}_{{\textbf{w}},S,\ell}$ is also injective.
-From 1.4 we have that, if the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective then it is also bi-Lipschitz. Let $a\leq b$ denote its bi-Lipschitz constants.
+From ˜1.4 we have that, if the map $\hat{\Phi}_{{\textbf{w}},S}$ is injective then it is also bi-Lipschitz. Let $a\leq b$ denote its bi-Lipschitz constants.
 Compositions of two Lipschitz maps is Lipschitz, hence $\Psi_{{\textbf{w}},S,\ell}$ is Lipschitz.
 Furthermore, an upper Lipschitz constant of $\Psi_{{\textbf{w}},S,\ell}$ is $\lVert\ell\rVert b$, where $\lVert\ell\rVert=\sigma_{1}(\ell)$ is the largest singular value of $\ell$.
-Finally from 3.2 and 3.3 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, $T=(N!)^{2p}$, $F_{a}=Ran(L_{\pi_{1},\dots,\pi_{2p}})$ we have that for a generic linear map $\ell$, for all $x,y\in{\mathscr{V}}$,
+Finally from ˜3.2 and ˜3.3 with $r=m-2d+d_{G}$, $s=2d-d_{G}$, $T=(N!)^{2p}$, $F_{a}=Ran(L_{\pi_{1},\dots,\pi_{2p}})$ we have that for a generic linear map $\ell$, for all $x,y\in{\mathscr{V}}$,
 $\lVert\Psi_{{\textbf{w}},S,\ell}(x)-\Psi_{{\textbf{w}},S,\ell}(y)\rVert$ $=\lVert\ell(D(x,y))\rVert\geq$
@@
 In previous sections we constructed the embedding $\hat{\Phi}_{{\textbf{w}},S}$ of the quotient space $\hat{{\mathscr{V}}}$ into an Euclidean space $\mathbb{R}^{m}$.
-In this section we prove 1.7 and 1.8 that show that when $\hat{\Phi}_{{\textbf{w}},S}$ is injective, every continous or Lipschitz map $f:\hat{{\mathscr{V}}}\to H$ factors through $\hat{\Phi}_{{\textbf{w}},S}$.
-The proof of 1.7 is based on Kirszbraun extension theorem [kirszbraun1934zusammenziehende] which is re-stated here for the reader’s convenience:
-Theorem 4.1 (Kirszbraun extension theorem[kirszbraun1934zusammenziehende]).
+In this section we prove ˜1.7 and ˜1.8 that show that when $\hat{\Phi}_{{\textbf{w}},S}$ is injective, every continous or Lipschitz map $f:\hat{{\mathscr{V}}}\to H$ factors through $\hat{\Phi}_{{\textbf{w}},S}$.
+The proof of ˜1.7 is based on Kirszbraun extension theorem [27] which is re-stated here for the reader’s convenience:
+Theorem 4.1 (Kirszbraun extension theorem[27]).
 Let $E\subset H_{1}$ be an arbitrary subset of a Hilbert space $H_{1}$ and $f:E\mapsto H_{2}$ be a Lipschitz function to another Hilbert space $H_{2}$.
 Then there exists an extension $F:H_{1}\rightarrow H_{2}$ of $f$ to the entire space $H_{1}$ that has the same Lipschitz constant as the original function $f$.
-Proof of 1.7.
+Proof of ˜1.7.
   1. 1.
@@
     $\leq\operatorname{Lip}(T)a\lVert x-y\rVert.$
-The second universality result 1.8 applies to the class of continuous functions (instead of Lipschitz functions).
-The proof of 1.8 follows from the following extension of Tietze’s theorem [urysohn1925machtigkeit, dugundji1951extension].
-Theorem 4.2 (Dugundji-Tietze [dugundji1951extension]).
+The second universality result ˜1.8 applies to the class of continuous functions (instead of Lipschitz functions).
+The proof of ˜1.8 follows from the following extension of Tietze’s theorem [35, 21].
+Theorem 4.2 (Dugundji-Tietze [21]).
 Let $X$ be a metric space and $A$ a closed subset of $X$.
@@
 Given $f:A\mapsto L$ a continuous map, there exists a continuous extension $F:X\mapsto L$ such as $F(X)$ is a subset of the convex hull of $f(A)$.
-Proof of 1.8.
+Proof of ˜1.8.
 Let $S=\Phi_{{\textbf{w}},S}({\mathscr{V}})\subset\mathbb{R}^{m}$.
@@
 Note $t$ is continuous since $\hat{\Phi}_{{\textbf{w}},S}$ is bi-Lipschitz on $\hat{{\mathscr{V}}}$.
-By 4.2 there exists a continuous extension $T:\mathbb{R}^{m}\rightarrow L$ of $t$ that satisfies the convex hull property.
-[…]
+By ˜4.2 there exists a continuous extension $T:\mathbb{R}^{m}\rightarrow L$ of $t$ that satisfies the convex hull property.